### v1

# Conformal Prediction in Hierarchical Classification

Abstract
Conformal prediction has emerged as a widely used framework for constructing valid prediction sets in classification and regression tasks. In this work, we extend the split conformal prediction framework to hierarchical classification, where prediction sets are commonly restricted to internal nodes of a predefined hierarchy, and propose two computationally efficient inference algorithms. The first algorithm returns internal nodes as prediction sets, while the second relaxes this restriction, using the notion of representation complexity, yielding a more general and combinatorial inference problem, but smaller set sizes. Empirical evaluations on several benchmark datasets demonstrate the effectiveness of the proposed algorithms in achieving nominal coverage.
Machine Learning, ICML
[…]
1 Introduction
In multi-class classification, a classifier can be uncertain about the predicted class label for a given test instance. In such cases, it can be beneficial to return set-valued predictions, i.e. sets of classes rather than individual classes. This is particularly relevant in hierarchical classification, where the class space is organized in a hierarchical structure, such as in medical diagnosis, where diseases are organized in a tree structure based on the International Classification of Diseases (ICD) (World Health Organization, 1978). In hierarchical classification, set-valued predictions are often limited to specific internal nodes in the hierarchy. Such sets have a clear semantic interpretation and can be constructed using efficient inference algorithms (Alex Freitas, 2007; Bi & Kwok, 2015; Rangwala & Naik, 2017; Yang et al., 2017; Wang et al., 2021; Valmadre, 2022). However, restricting predictions in this way can affect performance. For this reason, some authors allow any subset of classes as a prediction in hierarchical classification (Oh, 2017). Recently, a set-valued prediction framework for hierarchical classification that makes a compromise between the two aforementioned extremes has been proposed (Mortier et al., 2022). Compared to approaches that predict a single node of the hierarchy, this framework allows the user to restrict the representation complexity of a predicted set. The main idea is to return a restricted number of internal nodes of the hierarchy as candidate sets instead of a single node. However, this framework lacks a formal statistical guarantee and relies on well-calibrated probabilistic models.
Contributions. In this work, we build further upon the concept of representation complexity and extend the split conformal prediction framework to the hierarchical multi-class classification setting. Conformal prediction is a framework that provides valid and efficient set-valued predictions for any learning algorithm, without making any assumptions on the underlying data distribution (Vovk et al., 2005). In particular, given a trained (hierarchical) classifier, a desired coverage level $1-\alpha\in(0,1)$, $N$ calibration samples $\{(\bm{x}_{i},y_{i})\}_{i=1}^{N}$ and a test sample $(\bm{x}_{N+1},y_{N+1})$, drawn i.i.d. from an unknown distribution $P$, we would like to construct a prediction set $\hat{Y}\in 2^{\mathcal{Y}}$ with representation complexity $r$ for the test instance $(\bm{x}_{N+1},y_{N+1})$, such that the following marginal validity guarantee holds:
$\mathbb{P}\left[y_{N+1}\in\hat{Y}(\bm{x}_{N+1})\right]\geq 1-\alpha\,,\quad R_{\mathcal{T}}(\hat{Y})\leq r\,,$ (1)
where $R_{\mathcal{T}}(\hat{Y})$ denotes the representation complexity of the set $\hat{Y}$, given some tree structure $\mathcal{T}$. The probability above is taken over all $N+1$ samples and we require that (1) holds for any fixed $\alpha,N,r$ and $P$.
After reviewing some essentials on hierarchical probabilistic classification and conformal prediction in Section 2, we propose in Section 3 two algorithms that construct valid prediction sets, in the sense of satisfying (1). The first algorithm is designed for the restrictive case $r=1$, while the second extends to $r>1$. Moreover, we show that our algorithms possess distribution-free finite sample guarantees. Finally, in Section 4, we illustrate the notion of representation complexity and evaluate the proposed algorithms in terms of coverage and efficiency on a range of benchmark datasets.
$v_{1}=\{1,2,3,4,5,6,7,8\}$$v_{2}=\{1,2,3,4\}$$v_{4}=\{1,2\}$ $\begin{subarray}{c}v_{8}=\{1\},\\ 0.15\end{subarray}$ $\begin{subarray}{c}v_{9}=\{2\},\\ 0.13\end{subarray}$ $v_{5}=\{3,4\}$ $\begin{subarray}{c}v_{10}=\{3\},\\ 0.08\end{subarray}$ $\begin{subarray}{c}v_{11}=\{4\},\\ 0.125\end{subarray}$ $v_{3}=\{5,6,7,8\}$$v_{6}=\{5,6\}$ $\begin{subarray}{c}v_{12}=\{5\},\\ 0.14\end{subarray}$ $\begin{subarray}{c}v_{13}=\{6\},\\ 0.125\end{subarray}$ $v_{7}=\{7,8\}$ $\begin{subarray}{c}v_{14}=\{7\},\\ 0.125\end{subarray}$ $\begin{subarray}{c}v_{15}=\{8\},\\ 0.125\end{subarray}$
Figure 1: An example of a tree structure $\mathcal{T}$ with class space $\mathcal{Y}=\{1,\ldots,8\}$ and nodes $\mathcal{V}_{\mathcal{T}}=\{v_{1},\ldots,v_{15}\}$. The root $v_{1}$ represents the class space $\mathcal{Y}$ and leaves $\{v_{8},\ldots,v_{15}\}$ represent the individual classes. The numbers in the leaf nodes represent the class probabilities for some instance $\bm{x}$.
2 Background and related work
2.1 Hierarchical multi-class classification
In a standard multi-class classification setting one assumes that training and test data are i.i.d. according to an unknown distribution $P(\bm{x},y)$ on $\mathcal{X}\times\mathcal{Y}$, with $\mathcal{X}$ some instance space (e.g. images, documents, etc.) and $\mathcal{Y}=\{c_{1},\ldots,c_{K}\}$ a class space consisting of $K$ classes. Probabilistic multi-class classifiers estimate the conditional class probabilities $P(\cdot\,|\,\bm{x})$ over $\mathcal{Y}$, with properties $\forall c\in\mathcal{Y}:0\leq P(c\,|\,\bm{x})\leq 1\,,\sum_{c\in\mathcal{Y}}P(c\,|\,\bm{x})=1\,.$ This distribution can be estimated using a wide range of well-known probabilistic models, such as logistic regression, linear discriminant analysis, gradient boosting trees or neural networks with a softmax output layer. At prediction time, set-valued prediction algorithms, such as split conformal prediction, return sets $\hat{Y}$ that are subsets of $\mathcal{Y}$. The probability mass of such a set can be computed as $P(\hat{Y}\,|\,\bm{x})=\sum_{c\in\hat{Y}}P(c\,|\,\bm{x})$.
However, in this paper we will consider a hierarchical multi-class classification setting. Hence, we assume that a domain expert has defined a hierarchy over the class space, in the form of a tree structure $\mathcal{T}$ that contains in general $M$ nodes. $\mathcal{V}_{\mathcal{T}}=\{v_{1},\ldots,v_{M}\}$ will denote the set of nodes and every node identifies a set of classes. As special cases, the root $v_{1}$ represents the class space $\mathcal{Y}$, and the leaves represent individual classes – see Figure 1 for a simple example. In hierarchical classification, one typically makes the strong restriction $\hat{Y}\in\mathcal{V}_{\mathcal{T}}$ for predicted sets – see e.g. Bi & Kwok (2015). The probability mass $P(v\,|\,\bm{x})$ of such a set can be computed using the chain rule of probability:
$P(v\,|\,\bm{x})=\prod_{v^{\prime}\in\mathrm{Path}(v)}P(v^{\prime}\,|\,\mathrm{pa}(v^{\prime}),\bm{x})\,,$ (2)
where $\mathrm{Path}(v)$ is a set of nodes on the path connecting the node $v$ and the root of the tree structure. $\mathrm{pa}(v)$ gives the parent of node $v$ and $P(v\,|\,\mathrm{pa}(v),\bm{x})$ represents the branch probability of node $v$ given its parent $\mathrm{pa}(v)$. Note that for the root node $v_{1}$ one has $P(v_{1}\,|\,\mathrm{pa}(v_{1}),\bm{x})=1$. In Figure 1, the branch probabilities of the root node $v_{1}$ are given by $P(v_{2}\,|\,v_{1},\bm{x})=0.485$ and $P(v_{3}\,|\,v_{1},\bm{x})=0.515$. In order to estimate the branch probabilities, one can train any multi-class probabilistic classifier in each internal node of the tree. Classical models of that kind include nested dichotomies (Fox, 1997; Frank & Kramer, 2004; Melnikov & Hüllermeier, 2018), conditional probability estimation trees (Beygelzimer et al., 2009) and probabilistic classifier trees (Dembczyński et al., 2016). In neural networks with a hierarchical softmax output layer, all nodes are trained simultaneously (Morin & Bengio, 2005).
2.2 Inference in hierarchical classification
In this work, we do not focus on the training algorithms. Instead, we assume that a probabilistic model $\hat{P}$ has been estimated, either with classical models or using a hierarchical factorization as in (2), and we focus on the prediction task. In particular, we would like to predict sets $\hat{Y}\in 2^{\mathcal{Y}}$ that satisfy (1), with a restriction on the representation complexity of the predicted set (Mortier et al., 2022). The representation complexity is defined as the minimal number of nodes needed to represent the set $\hat{Y}$ in the tree structure $\mathcal{T}$. More formally, let $\mathcal{S}_{\mathcal{T}}(\hat{Y})$ denote the set of all disjoint combinations of tree nodes that represent the set $\hat{Y}$:
$\mathcal{S}_{\mathcal{T}}(\hat{Y})=\left\{\hat{V}\subset\mathcal{V}_{\mathcal{T}}:\bigcup_{v_{i}\in\hat{V}}v_{i}=\hat{Y}\land\bigcap_{v_{i}\in\hat{V}}v_{i}=\emptyset\right\}\,.$
Then, we define the representation complexity of the prediction $\hat{Y}$ as
$R_{\mathcal{T}}(\hat{Y})=\min_{\hat{V}\in\mathcal{S}_{\mathcal{T}}(\hat{Y})}|\hat{V}|\,,$ (3)
with $|\hat{V}|$ the cardinality of $\hat{V}$. For example, the set $\hat{Y}=\{1,2,4,7,8\}$ of the hierarchy shown in Figure 1 has representation complexity three, because three nodes are needed to represent this set: $v_{4}$, $v_{7}$ and $v_{11}$.
2.3 Randomized nested prediction sets with split conformal prediction
Now we will review a general procedure for valid set-valued predictions in flat classification (i.e. ignoring the hierarchy), following the work of Gupta et al. (2022); Romano et al. (2020); Angelopoulos et al. (2020). Compared to traditional conformal prediction—which starts from the notion of a nonconformity score—this procedure departs from a sequence of nested set-valued predictions, where the set size depends on a threshold $\tau$. Furthermore, an independent calibration set is used to tune the threshold $\tau$ such that $(1-\alpha)$-coverage is guaranteed on future test samples. The use of a single calibration set is better known as split conformal prediction and gives rise to computationally efficient valid set-valued predictions (Papadopoulos et al., 2002; Lei et al., 2018).
More formally, let $\{(\bm{x}_{i},y_{i})\}_{i=1}^{N}$ be an i.i.d. sequence of $N$ samples from the unknown distribution $P$, and assume that these samples were not used for model training. Let $\hat{Y}(\bm{x},u,\tau):\mathcal{X}\times[0,1]\times\mathbb{R}\rightarrow 2^{\mathcal{Y}}$ be a set-valued predictor. In the spirit of Romano et al. (2020); Angelopoulos et al. (2020), the second argument $u$ represents a random draw from a uniform distribution $\mathcal{U}(0,1)$ and is included to allow for randomized prediction sets. The third argument $\tau$ is a threshold that controls the size of the predicted set. Assume that the sets are indexed by $\tau$ and are nested:
$\hat{Y}(\bm{x},u,\tau_{1})\subseteq\hat{Y}(\bm{x},u,\tau_{2})\qquad\text{if}\quad\tau_{1}\leq\tau_{2}\,.$ (4)
In order to find the optimal threshold $\tau^{*}$ that guarantees $(1-\alpha)$-coverage, one needs to find the smallest $\tau$ such that the predicted set $\hat{Y}(\bm{x},u,\tau)$ contains at least $\lceil(N+1)(1-\alpha)\rceil$ samples:
$\tau^{*}=\inf\big{\{}\tau\in[0,1]:|\{i:y_{i}\in\hat{Y}(\bm{x}_{i},u_{i},\tau)\}|$
$\geq\lceil(1-\alpha)(N+1)\rceil\big{\}}\,.$ (5)
The set-valued predictor $\hat{Y}(\bm{x}_{N+1},u_{N+1},\tau^{*})$, with $\tau^{*}$ in (2.3), has marginal validity guarantees, as shown by the theorem below.
Theorem 2.1 (Marginal validity of nested conformal prediction (Angelopoulos et al., 2020)).
Assume an exchangeable sequence $\{(\bm{x}_{i},y_{i},u_{i})\}_{i=1}^{N+1}$ and let $\hat{Y}(\bm{x},u,\tau)$ be a set-valued predictor that satisfies (4). Furthermore, assume that $\exists\tau\in\mathbb{R}:\hat{Y}(\bm{x},u,\tau)=\mathcal{Y}$. Then, for $\tau^{*}$ in (2.3) and any $\alpha\in(0,1)$, the following marginal coverage guarantee holds:
$1-\alpha\leq\mathbb{P}\big{[}y_{N+1}\in\hat{Y}(\bm{x}_{n+1},u_{N+1},\tau^{*})\big{]}\,.$
Moreover, if the sets grow smoothly in $\tau$, then the following upper bound holds:
$\mathbb{P}\big{[}y_{N+1}\in\hat{Y}(\bm{x}_{n+1},u_{N+1},\tau^{*})\big{]}\leq 1-\alpha+\frac{1}{N+1}\,.$
In the literature, several conformal prediction methods can be found that satisfy (4). For example, in Romano et al. (2020), the following set-valued predictor, called adaptive prediction sets (APS), is proposed:
$\hat{Y}_{APS}(\bm{x},u,\tau)=\big{\{}y\in\mathcal{Y}:\hat{\rho}(y;\bm{x})+u\cdot\hat{P}(y\,|\,\bm{x})\leq\tau\big{\}}\,,$ (6)
with $\hat{\rho}(y;\bm{x})=\sum_{y^{\prime}\in\mathcal{Y}}\hat{P}(y^{\prime}\,|\,\bm{x})\mathbbm{1}_{\hat{P}(y^{\prime}\,|\,\bm{x})>\hat{P}(y\,|\,\bm{x})}$ the probability mass of the labels more likely than $y$. At the heart of their method is the randomization term $u\cdot\hat{P}(y\,|\,\bm{x})$, which has been introduced in order to achieve exact nominal coverage. In addition, by means of the cumulative distribution $\hat{\rho}(y;\bm{x})$, this method allows to adapt effectively to complex data distributions, achieving a better conditional coverage compared to alternative approaches that rely on the mode of the distribution, such as the least ambiguous classifier proposed by Sadinle et al. (2019). In Angelopoulos et al. (2020), a variant of the above method has been proposed, called the regularized adaptive prediction sets (RAPS) method. By introducing a regularization term in (6), which needs to be tuned by means of additional tuning data, small set sizes are encouraged, in particular, when faced with noisy probability estimates for classes with low probability.
Alternative conformal prediction methods have been introduced, focusing on minimizing set size (Sadinle et al., 2019) and achieving approximate coverage across the entire feature space (Foygel Barber et al., 2021; Cauchois et al., 2021; Romano et al., 2019). In Angelopoulos et al. (2023), a conformal prediction method is proposed that controls the risk with respect to any bounded non-increasing loss function. Interestingly, this method is also applicable to the setting of hierarchical classification, i.e. through the use of the tree-distance loss, introduced in Bi & Kwok (2015) as a way of evaluating set-valued predictions in multi-label classification. However, set-valued predictions of that kind lack meaningful and practical interpretation, compared to methods that rely on the traditional notion of coverage. Furthermore, this method is not directly applicable to our work, since it does not allow the incorporation of representation complexity, given the properties of the loss function above.
3 Proposed methods
Building upon the concepts in Section 2.3, we are now ready to discuss two algorithms that provide valid set-valued predictions for hierarchical classification that satisfy (1) above. The first algorithm provides marginal validity guarantees in classical hierarchical classification settings, i.e. where sets are restricted to nodes of the tree structure, thus $R_{\mathcal{T}}(\hat{Y})=1$. The second algorithm provides marginal validity guarantees in restricted set-valued prediction settings, i.e., where the representation complexity of the predicted set is restricted to $R_{\mathcal{T}}(\hat{Y})\leq r$ for a user-defined value $r$. Note that traditional conformal prediction in flat classification is obtained for the second algorithm when there is no restriction on the representation complexity (e.g. $r=K$).
3.1 Conformal restricted set-valued prediction (CRSVP)
The first approach, called conformal restricted set-valued prediction (CRSVP), predicts sets that are restricted to nodes of the hierarchy, thus having representation complexity $R_{\mathcal{T}}(\hat{Y})=1$. Assume one has fitted a hierarchical classifier using a training dataset, as described in Section 2.1 and let $\hat{y}(\bm{x})$ denote the mode of the distribution $\hat{P}(.\,|\,\bm{x})$, i.e. the leaf node with highest probability mass. For our first approach, we consider the following set-valued predictor:
$\hat{Y}_{1}(\bm{x},u,\tau)=\operatorname*{\arg\max}_{\hat{Y}\in\mathrm{Path}(\hat{y}(\bm{x}))}\big{\{}|\hat{Y}|:\hat{P}(\hat{Y}\,|\,\bm{x})+$
$u\cdot\hat{P}(\mathrm{pa}(\hat{Y})\setminus\hat{Y}\,|\,\bm{x})\leq\tau\big{\}}\,,$ (7)
with $\mathrm{Path}(v)$ the path to the root, and $\mathrm{pa}(v)$ the parent of a node $v$, as defined in Section 2. $\hat{P}(\hat{Y}\,|\,\bm{x})$ can be computed using (2), but also observe that the probability mass of an internal node $v$ corresponds to $\hat{P}(v\,|\,\bm{x})=\sum_{c\in v}\hat{P}(c\,|\,\bm{x})$, i.e., the sum of the probability masses of the leaf nodes that are descendants of $v$. However, computing the probability mass of an internal node and its associated set of classes in this way is computationally less efficient than using the chain rule in (2), when hierarchical classifiers have been fitted during training.
It is clear that the set-valued predictor in (3.1) satisfies the nested property in (4). Indeed, starting from the mode of the distribution $\hat{y}(\bm{x})$, every node on the path connecting that node and the root is nested by definition of the hierarchical tree structure $\mathcal{T}$. In line with Angelopoulos et al. (2020), the first term increases as we move up the tree, while the second term contains a random draw from the uniform distribution $\mathcal{U}(0,1)$, for handling discrete jumps in probability mass when following the path towards the root. The randomization is needed to prevent over-coverage.
The algorithm for calculating the threshold in (2.3) using the calibration dataset is presented in Algorithm 1. This algorithm first computes the mode of the estimated class probabilities for each instance in the calibration dataset. For small hierarchies, this can be done using exhaustive search, but more efficient inference algorithms have been developed when the computational cost becomes a burden (Dembczyński et al., 2012; Kumar et al., 2013; Mena Waldo et al., 2015). Subsequently, starting from the mode, Algorithm 1 computes for every calibration instance the internal node, on the path to the root, that also includes the true class label. Nonconformity scores that incorporate a randomization component are constructed for these internal nodes, and the final threshold is deduced from the scores. Algorithm 1 has a worst case $\mathcal{O}(N\,K)$ time complexity when exhaustive search is applied in the first step.
The pseudocode for computing the set-valued prediction for a new test instance, as defined in (3.1), is shown in Algorithm 2. Given a hierarchical classifier and a balanced hierarchy, the computational complexity during test time is given by $\mathcal{O}(\log K)$.
Algorithm 1 CRSVP calibration – Input: $\{(\bm{x}_{i},y_{i},u_{i})\}_{i=1}^{N},\hat{P},\mathcal{V}_{\mathcal{T}}$, Output: Threshold in (2.3).
1:  for $i=1,\ldots,N$ do
2:    $\hat{Y}\leftarrow\operatorname*{\arg\max}_{c\in\mathcal{Y}}\hat{P}(c\,|\,\bm{x}_{i})$
3:    $\hat{p}_{\hat{Y}}\leftarrow\hat{P}(\hat{Y}\,|\,\bm{x}_{i}),\hat{p}_{\hat{Y}^{\prime}}\leftarrow 0$
4:    while $y_{i}\notin\hat{Y}$ do
5:       $\hat{p}_{\hat{Y}^{\prime}}\leftarrow\hat{p}_{\hat{Y}}$
6:       $\hat{Y}\leftarrow\mathrm{pa}(\hat{Y}),\hat{p}_{\hat{Y}}\leftarrow\hat{P}(\hat{Y}\,|\,\bm{x}_{i})$
7:    end while
8:    $\tau_{i}\leftarrow\hat{p}_{\hat{Y}}-u_{i}\cdot(\hat{p}_{\hat{Y}}-\hat{p}_{\hat{Y}^{\prime}})$
9:  end for
10:  $\tau^{*}\leftarrow$ the $\lceil(1-\alpha)(N+1)\rceil$-th largest value in $\{\tau_{i}\}_{i=1}^{N}$
11:  return $\tau^{*}$
Algorithm 2 CRSVP inference – Input: $\bm{x},\tau^{*},u,\hat{P},\mathcal{V}_{\mathcal{T}}$, Output: Set-valued prediction in (3.1).
1:  $\hat{Y}\leftarrow\operatorname*{\arg\max}_{c\in\mathcal{Y}}\hat{P}(c\,|\,\bm{x}),\hat{Y}^{\prime}\leftarrow\emptyset$
2:  $\hat{p}_{\hat{Y}}\leftarrow\hat{P}(\hat{Y}\,|\,\bm{x}),\hat{p}_{\hat{Y}^{\prime}}\leftarrow 0$
3:  while $|\hat{Y}^{\prime}|\neq K$ do
4:    if $\hat{p}_{\hat{Y}}+u\cdot\hat{P}(\mathrm{pa}(\hat{Y})\setminus\hat{Y}\,|\,\bm{x})>\tau^{*}$ then
5:       break
6:    end if
7:    $\hat{Y}^{\prime}\leftarrow\hat{Y},\hat{Y}\leftarrow\mathrm{pa}(\hat{Y})$
8:    $\hat{p}_{\hat{Y}^{\prime}}\leftarrow\hat{p}_{\hat{Y}},\hat{p}_{\hat{Y}}\leftarrow\hat{P}(\hat{Y}\,|\,\bm{x})$
9:  end while
10:  return $\hat{Y}^{\prime}$
3.2 Conformal set-valued prediction with representation complexity (CRSVP-$r$)
It should be obvious for the reader that sets of representation complexity one quickly become very big, since they correspond to internal nodes of the hierarchy that exhibit a predefined coverage level. Therefore, we present a second approach, called conformal set-valued prediction with representation complexity (CRSVP-$r$), that relaxes the representation complexity constraint to $R_{\mathcal{T}}(\hat{Y})\leq r$ with $r$ a user-defined parameter. Let
$\hat{\omega}(y;\bm{x})=\left\{y^{\prime}\in\mathcal{Y}:\hat{P}(y^{\prime}\,|\,\bm{x})\geq\hat{P}(y\,|\,\bm{x})\right\}$ (8)
be the set of classes that are more likely than the label $y$ for $\bm{x}$. As a crucial concept for the second approach, we first define the following optimization problem:
$A_{r}(y;\bm{x})=\operatorname*{\arg\min}_{\begin{subarray}{c}\hat{Y}\in 2^{\mathcal{Y}}:R_{\mathcal{T}}(\hat{Y})\leq r,\\ \hat{\omega}(y;\,\bm{x})\subseteq\hat{Y}\end{subarray}}|\hat{Y}|-\hat{P}(\hat{Y}\,|\,\bm{x})\,.$ (9)
We will refer to this optimization problem as finding the set of common ancestors. It can be thought of as a generalization of the well-known lowest common ancestor problem (Aho et al., 1973; Harel & Tarjan, 1984; Bender & Farach-Colton, 2000; Dash et al., 2013), i.e. instead of considering a single common ancestor of a set of leaves in a tree, we are looking for a set of $r$ non-overlapping ancestors whose descendants form a set of representation complexity $r$. The second term in the objective function is usually not considered in the lowest common ancestor problem, but it is needed here to make sure that the solution of (9) is unique. Since probabilities are bounded between zero and one, the second term only plays a role for sets that have the same cardinality. As an example, consider the hierarchy depicted in Figure 1 above. For $y=2$ we find $\hat{\omega}(y;\bm{x})=\{1,2,5\}$. The lowest common ancestor of this set is $A_{1}(y;\bm{x})=\{1,2,3,4,5,6,7,8\}$. The lowest set of common ancestors with representation complexity two is $A_{2}(y;\bm{x})=\{1,2,5\}$.
Let $\hat{Y}^{(1)},\ldots,\hat{Y}^{(K)}$ denote the ordered sequence of unique lowest common ancestors $A_{r}(y^{(1)};\bm{x}),\ldots,A_{r}(y^{(K)};\bm{x})$, with $y^{(1)},\ldots,y^{(K)}$ the classes sorted in decreasing order of probability for an instance $\bm{x}$. We use the ordered sequence of unique lowest common ancestors to introduce the following set-valued predictor:
$\hat{Y}_{\leq r}(\bm{x},u,\tau)=\operatorname*{\arg\max}_{\hat{Y}^{(k)}\in\{\hat{Y}^{(1)},\ldots,\hat{Y}^{(K)}\}}\big{\{}|\hat{Y}^{(k)}|:$
$\hat{P}(\hat{Y}^{(k-1)}\,|\,\bm{x})+u\cdot\hat{P}(\hat{Y}^{(k)}\setminus\hat{Y}^{(k-1)}\,|\,\bm{x})\leq\tau\big{\}}\,.$ (10)
Before discussing the set-valued predictor in (3.2), we first discuss the algorithmic aspects of computing the set of lowest common ancestors in (9), because the algorithmic insights are needed to formally prove that the sets returned by (3.2) are nested. At first impression, the combinatorial optimization problem in (9) looks challenging, but it can be solved in an efficient manner with a divide-and-conquer algorithm that is described in Algorithm S6. To this end, some data structures need to be initialized first in Algorithm S5. Both algorithms can be found in the supplementary materials.
For every internal node of the hierarchy, the optimization problem can be broken down into simpler subproblems of the same type. As an example, consider an internal node $v^{\prime}$ that has four children $(v^{\prime}_{1},v^{\prime}_{2},v^{\prime}_{3},v^{\prime}_{4})$ in a predefined hierarchy, and assume that we aim to find set of lowest common ancestors with representation complexity three for a set of classes $\hat{\omega}(y;\bm{x})$ that are all descendants of $v^{\prime}$. As potential divisions of the representation complexity, one could find one common ancestor as descendants of children $v^{\prime}_{1}$, $v^{\prime}_{3}$, and $v^{\prime}_{4}$, or as descendants of children $v^{\prime}_{1}$, $v^{\prime}_{2}$, and $v^{\prime}_{4}$, or as two descendants of $v^{\prime}_{1}$ combined with one descendant of $v^{\prime}_{4}$, etc. In fact, many combinations are possible. Speaking in formal combinatorial terminology, one has to consider for the node $v$ all weak compositions of the integer three into four elements. For each of these weak compositions, the optimization problem can be recursively divided into smaller problems with lower representation complexity, where the node $v$ is replaced by each of its non-zero children in the composition. However, implementing such a strategy in a recursive manner would heavily blow up the computations, since each of the smaller subproblems would need to be solved many times.
A dynamic programming implementation, which avoids recursion by solving the smaller subproblems in a bottom-up manner, before tackling the more challenging optimization problems higher up in the hierarchy, is able to solve (9) in an efficient manner for small values of $r$. The pseudocode of such an implementation is described in Algorithm S5 and S6. For every internal node, $r$ local optimization problems are solved by varying the local representation complexity $r_{i}$ from one to $r$, and the solutions of these optimization problems are stored. By visiting children before their parents get analysed, one can guarantee that all needed quantities are known when the different weak compositions in an internal node need to be investigated. The most critical step in the algorithm is line fifteen, i.e. where all compositions of the current representation complexity $r_{i}$ into $|T|$ elements are considered. Strictly speaking, this step has a runtime that is exponential in $r$. However, in practical situations, one would only be interested in sets with a representation complexity lower than five, so Algorithm S6 in general computes the exact solution to (9) in an efficient manner.
A similar reasoning can be made to prove that the sets returned by (3.2) are nested.
Theorem 3.1 (Nestedness of CRSVP-$r$).
Consider two threshold $\tau_{1}$ and $\tau_{2}$, with $\tau_{1}\leq\tau_{2}$, then $\hat{Y}_{\leq r}(\bm{x},u,\tau_{1})\subseteq\hat{Y}_{\leq r}(\bm{x},u,\tau_{2})$.
Proof.
Changing the threshold from $\tau_{1}$ to $\tau_{2}$ can have three consequences: (1) no additional class label can be added to $\hat{Y}_{\leq r}(\bm{x},u,\tau_{1})$, (2) the next class label from the list with sorted class probabilities is added to $\hat{Y}_{\leq r}(\bm{x},u,\tau_{1})$, or (3) more than one class label from the list with sorted class probabilities is added to $\hat{Y}_{\leq r}(\bm{x},u,\tau_{1})$. We only have to further investigate the second situation: the set-valued prediction does not change in the first situation, whereas the third situation can be reduced to the second situation by considering intermediate changes in threshold, so that only a single class is added at the same time.
For the second situation, assume that the maximally allowed representation complexity is $r$ and recall the divide-and-conquer strategy of Algorithm S6. We have to consider three subcases.
Case 1: Adding the new class $y^{\prime}$ does not change the solution of (9) found for $\tau_{1}$. This case does not need further attention, because this situation cannot occur. With the nonconformity score that is considered, the new class would have been added already with the threshold $s_{1}$, leading to a contradiction.
Case 2: Adding the new class $y^{\prime}$ changes the solution of (9) found for $\tau_{1}$, but that solution had a representation complexity that was strictly lower than the maximally allowed representation complexity $r$. The solution to the new optimization problem is the solution to the old optimization, appended with the leaf node $y^{\prime}$.
Case 3: Adding the new class $y^{\prime}$ changes the solution of (9) found for $\tau_{1}$, and that solution had a representation complexity that was equal to the maximally allowed representation complexity $r$. Adding the leaf node $y^{\prime}$ to the solution obtained for $\tau_{1}$ is in this case not possible, since this would lead to a violation of the representation complexity constraint. The new solution will be different. Let’s think about what changes are possible compared to the solution for $\tau_{1}$. When considering the dynamic programming strategy of Algorithm S6, one can observe that only a different decision can be taken in internal nodes on the path from the leaf $y^{\prime}$ to the root. Moreover, a different decision will be taken in at least one internal node on that path. In contrast, no new decision can be taken on internal nodes that are not on the path from $y^{\prime}$ to the root, because the information in these nodes has not changed. So, let’s analyze what changes are possible in the internal nodes on the path from $y^{\prime}$ to the root. In such an internal node, it can happen that the optimal weak composition for $\tau_{1}$ will change, but it can only change in such a way that the branch leading to $y^{\prime}$ is increased with one unit. As a result, one of the other branches has to decrease with one unit. Two descendants of that branch that were part of the optimal solution will have to be merged, resulting in a bigger set that corresponds to the ancestor of these nodes. Likewise, in the branch that contains $y^{\prime}$, the set size will also increase, because the new leaf $y^{\prime}$ needs to incorporated. Thus, for every internal node along the path from $y^{\prime}$ to the root, it happens that the optimal sets are nested when increasing the threshold. ∎
The pseudocode for calculating the threshold in (2.3) and set-valued predictions in (3.2) are presented in Algorithm 3 and 4. Given a hierarchical classifier and a binary tree, the computational complexity during test time is given as follows. Initialization in Algorithm S5 is given by $\mathcal{O}(K\log K)$, since in the worst case, each node in the tree must be traversed. Furthermore, calculating the set of lowest common ancestors in Algorithm S6 is given by $\mathcal{O}(K2^{r}r)$. Note that in practice, $r$ is small and the computational complexity is therefore manageable. The overall computational complexity of Algorithm 4 is then given by $\mathcal{O}(lK\log K+lK2^{r}r)$, where $l$ is the number of unique lowest common ancestors.
Algorithm 3 CRSVP-$r$ calibration – Input: $\{(\bm{x}_{i},y_{i},u_{i})\}_{i=1}^{N},r,\hat{P},\mathcal{V}_{\mathcal{T}}$, Output: Threshold in (2.3).
1:  for $i=1,\ldots,N$ do
2:    $k\leftarrow 1$
3:    $\hat{Y}^{(k-1)}\leftarrow\emptyset,\hat{Y}^{(k)}\leftarrow\emptyset$
4:    $\hat{p}_{\hat{Y}^{(k-1)}}\leftarrow 0,\hat{p}_{\hat{Y}^{(k)}}\leftarrow 0$
5:    while $k\leq K$ do
6:       if $y_{i}^{(k)}\notin\hat{Y}^{(k-1)}$ then
7:          $\hat{Y}^{(k)^{\prime}}\leftarrow A_{r}(y^{(k)}_{i};\bm{x}_{i})$ (by means of Algorithm S5 and  S6)
8:          $\hat{p}_{\hat{Y}^{(k)^{\prime}}}\leftarrow\hat{P}(\hat{Y}^{(k)^{\prime}}\,|\,\bm{x}_{i})$
9:          if $|\hat{Y}^{(k)^{\prime}}|\neq|\hat{Y}^{(k-1)}|$ then
10:             $\hat{Y}^{(k)},\hat{p}_{\hat{Y}^{(k)}}\leftarrow\hat{Y}^{(k)^{\prime}},\hat{p}_{\hat{Y}^{(k)^{\prime}}}$
11:          end if
12:       end if
13:       if $y_{i}^{(k)}=y_{i}$ then
14:          break
15:       end if
16:       $k\leftarrow k+1$
17:    end while
18:    $\tau_{i}\leftarrow\hat{p}_{\hat{Y}^{(k)}}-u_{i}\cdot(\hat{p}_{\hat{Y}^{(k)}}-\hat{p}_{\hat{Y}^{(k-1)}})$
19:  end for
20:  $\tau^{*}\leftarrow$ the $\lceil(1-\alpha)(N+1)\rceil$-th largest value in $\{\tau_{i}\}_{i=1}^{N}$
21:  return $\tau^{*}$
Algorithm 4 CRSVP-$r$ inference – Input: $\bm{x},\tau^{*},u,r,\hat{P},\mathcal{V}_{\mathcal{T}}$, Output: Set-valued prediction in (3.2).
1:  $k\leftarrow 1$
2:  $\hat{Y}^{(k-1)}\leftarrow\emptyset,\hat{Y}^{(k)}\leftarrow\emptyset$
3:  $\hat{p}_{\hat{Y}^{(k-1)}}\leftarrow 0,\hat{p}_{\hat{Y}^{(k)}}\leftarrow 0$
4:  while $k\leq K$ do
5:    if $y^{(k)}\notin\hat{Y}^{(k-1)}$ then
6:       $\hat{Y}^{(k)^{\prime}}\leftarrow A_{r}(y^{(k)};\bm{x})$ (by means of Algorithm S5 and  S6)
7:       $\hat{p}_{\hat{Y}^{(k)^{\prime}}}\leftarrow\hat{P}(\hat{Y}^{(k)^{\prime}}\,|\,\bm{x})$
8:       if $|\hat{Y}^{(k)^{\prime}}|\neq|\hat{Y}^{(k-1)}|$ then
9:          $\hat{Y}^{(k)},\hat{p}_{\hat{Y}^{(k)}}\leftarrow\hat{Y}^{(k)^{\prime}},\hat{p}_{\hat{Y}^{(k)^{\prime}}}$
10:       end if
11:       if $\hat{p}_{\hat{Y}^{(k-1)}}+u\cdot(\hat{p}_{\hat{Y}^{(k)}}-\hat{p}_{\hat{Y}^{(k-1)}})>\tau^{*}$ then
12:          break
13:       end if
14:       $k\leftarrow k+1$
15:    end if
16:  end while
17:  return $\hat{Y}^{(k-1)}$
4 Experimental results
4.1 Experimental setup
In this section, two types of experiments are considered in light of the concepts that have been discussed in previous sections. The first experiment demonstrates the notion of representation complexity using an example from the PlantCLEF 2015 dataset (Göeau et al., 2015). In the second experiment, different set-valued predictors are compared for solving problem (1) across four benchmark datasets, focusing on coverage, efficiency, and representation complexity. Summary statistics for these datasets are presented in Table 1. For all datasets, a predefined hierarchy $\mathcal{T}$ is extracted from the provided hierarchical labels.
All dataset images are converted to RGB format with a resolution of $200\times 200$ pixels. Hidden representations are obtained using an EfficientNet convolutional neural network (Tan & Le, 2019) pretrained on ImageNet (Krizhevsky et al., 2017). These hidden representations are fed into a hierarchical softmax layer, with dropout set to 0.1. The cross-entropy loss is minimized using the Adam optimizer, with a learning rate of $1\times 10^{-5}$ and momentum set to 0.99. Training is performed with a batch size of 32 for up to 100 epochs, with early stopping applied after five epochs of no improvement. All models are trained end-to-end on a GPU using the PyTorch library (Paszke et al., 2019), with the following hardware configuration:
  • CPU: Intel i7-6800K 3.4 GHz (3.8 GHz Turbo Boost)
  • GPU: NVIDIA GTX 1080 Ti 11GB
  • RAM: 64GB DDR4-2666
All inference algorithms are implemented in C++ using the PyTorch C++ API (Paszke et al., 2019).
Table 1: Summary statistics and top-1 performance for all datasets. Notation: $K$ – number of classes, $N$ – number of samples, Top-1 acc. – top-1 accuracy of hierarchical classifier.
Dataset $K$ $N_{train}$ $N_{cal}$ $N_{test}$ Top-1 acc.
CIFAR-10 10 50000 5000 5000 0.8817
Caltech-101 97 4338 2169 2169 0.9039
Caltech-256 256 14890 7445 7445 0.7578
PlantCLEF 2015 1000 91758 10723 10723 0.4156
4.2 Illustrations
(a) Cirsium arvense (L.) Scop.
(b) Lotus corniculatus L.
Figure 2: Two example images corresponding to two species from the PlantCLEF 2015 dataset. Left: Cirsium arvense (L.) Scop., right: Lotus corniculatus L..
Table 2: Results for CIFAR-10, Caltech-101, Caltech-256 and PlantCLEF 2015. Coverage, efficiency, and representation complexity for the following unrestricted set-valued predictors: LAC, NPS, APS, and restricted set-valued predictors: NCRSVP, CRSVP, NCRSVP-$r$, and CRSVP-$r$. The confidence level is set to 90%, and calibration and test sets are resampled 20 times. Calculated standard deviations are negligible and have therefore been excluded from the table.
Alg. LAC NPS APS NCRSVP CRSVP NCRSVP-1 CRSVP-1 NCRSVP-3 CRSVP-3
Dataset
CIFAR-10 Cov. 0.8987 0.9998 0.9008 1.0 0.9019 1.0 0.9017 0.9998 0.9016
Size 1.0412 4.9756 1.2739 7.5735 1.8418 7.7111 1.7554 5.2632 1.284
Repr. comp. 1.0357 2.5707 1.2705 1.0 1.0 1.0 1.0 1.9872 1.249
Caltech-101 Cov. 0.9025 0.9984 0.9032 0.9992 0.9012 0.9993 0.9023 0.9980 0.9036
Size 0.9322 17.114 1.5461 48.2421 5.1427 49.9799 4.9962 33.3329 2.1821
Repr. comp. 1.0 11.0477 1.5955 1.0 1.0 1.0 1.0 1.7412 1.243
Caltech-256 Cov. 0.8992 0.9821 0.9017 0.9980 0.8995 0.9993 0.9067 0.9912 0.9079
Size 2.3446 28.0667 4.7988 210.3485 43.8557 232.469 44.6294 105.4433 17.7206
Repr. comp. 2.2832 19.6563 4.4863 1.0 1.0 1.0 1.0 1.82 1.7149
PlantCLEF 2015 Cov. 0.8991 0.9183 0.8988 1.0 0.9005 1.0 0.9085 1.0 0.9134
Size 73.5715 106.05 91.8655 1000.0 676.9661 1000.0 702.7522 1000.0 591.3467
Repr. comp. 60.2026 82.6895 72.8162 1.0 1.0 1.0 1.0 1.0 1.6537
To illustrate the notion of representation complexity, we consider the PlantCLEF 2015 dataset (Göeau et al., 2015). This dataset consists of over one hundred thousand images representing 1,000 species of trees, herbs, and ferns native to the Western European region. This dataset is characterized by significant class ambiguity, making accurate predictions on species level often impossible. From an uncertainty point of view, this challenge is particularly interesting, as even biological experts sometimes encounter uncertainty when validating visual observations of living organisms.
In Figure 2, two example images are shown, corresponding to two species from the PlantCLEF 2015 dataset. The left image shows Cirsium arvense (L.) Scop., a species of the Asteraceae family, commonly known as the creeping thistle. Two set-valued predictions by means of CRSVP-$r$ (Algorithm 4) for $r=1,3$ are given as follows:
[…]
With representation complexity restricted to one, the set-valued prediction corresponds to a genus level, encompassing six species. By increasing the representation complexity to three, the set-valued prediction becomes more precise, encompassing three species of the same genus, while including the true label.
The right image shows Lotus corniculatus L., a member of the Fabaceae family, commonly referred to as bird’s-foot trefoil. Similarly, we consider two set-valued predictions obtained by CRSVP-$r$ for $r=1$ and $r=3$:
$\hat{Y}_{\leq 1}(\bm{x},u,\tau^{*})$ $=\mathcal{Y}$
[…]
In this case, the set-valued prediction with representation complexity one is imprecise, as it includes all 1,000 species. As before, increasing the representation complexity to three improves precision, i.e. narrowing the prediction set to three (visually-related) species, while including the true label.
4.3 Benchmarking results
In a second experiment, we compare various set-valued predictors that solve problem (1) across the four benchmark datasets listed in Table 1: CIFAR-10 (Krizhevsky et al., 2010), Caltech-101 (Li et al., 2003), Caltech-256 (Griffin et al., 2007), and PlantCLEF 2015 (Göeau et al., 2015). Specifically, we evaluate the following set-valued predictors: CRSVP, CRSVP-1, and CRSVP-$3$. Additionally, we demonstrate the usefulness of randomized prediction sets by considering the following naive (N) set-valued predictors: NCRSVP, NCRSVP-1, and NCRSVP-3. These correspond to setting $u$ to zero when calculating the threshold in (2.3) and constructing the prediction sets in (3.1) and (3.2), respectively. Finally, we include results for three baseline methods that produce valid set-valued predictions in flat classification (i.e. ignoring the hierarchy): the least ambiguous classifier (LAC), as proposed by Sadinle et al. (2019); adaptive prediction sets (APS), as defined in (6); and naive prediction sets (NPS), which correspond to APS without randomization (i.e. setting $u$ to zero).
The results are obtained by training a neural network with a hierarchical softmax layer, as discussed in Section 4.1. During inference, we resample the calibration and test sets 20 times and use a confidence level of 90%. The results are summarized in Table 2. For each experiment, we report coverage, efficiency, and representation complexity. Coverage is defined as the proportion of samples for which the true class is included in the prediction set. Efficiency is defined as the average size of the set-valued prediction. The results clearly indicate that naive set-valued predictors fail to deliver prediction sets with exact coverage, thereby highlighting the importance of randomized prediction sets. Moreover, increasing the representation complexity generally improves efficiency, demonstrating its practical value. In extreme cases, when representation complexity is unrestricted, such as with LAC, NPS, and APS, optimal performance in terms of efficiency is observed. However, this comes at the cost of significantly increased representation complexity in the prediction sets, in particular for large $K$, which may not be practical when predictions need to adhere to a predefined hierarchy.
5 Conclusion
In this work, we extended the split conformal prediction framework to hierarchical classification by introducing two novel set-valued prediction algorithms. The first algorithm generates valid set-valued predictions restricted to single nodes within a predefined hierarchy. We argued that this restriction can be limiting in certain applications. To address this limitation, we introduced a second algorithm that relaxes this constraint by incorporating the notion of representation complexity. Empirical evaluations on multiple benchmark datasets demonstrate the effectiveness of the proposed algorithms in achieving exact nominal coverage. In addition to exploring alternative constraints on the prediction sets, another interesting direction for future research would be to generalize our proposed methods to more complex structures, such as directed acyclic graphs.
Impact statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Supplementary materials
A.1 Computing the set of lowest common ancestors for CRSVP-$r$
In this section, we present a dynamic programming solution to (9). The pseudocode is outlined in Algorithms S5 and S6. For each internal node, $r$ local optimization problems are solved by varying the local representation complexity $r_{i}$ from 1 to $r$ and their solutions are stored. Processing children before parents ensures all necessary values are available when evaluating weak compositions within an internal node. The critical step, line 15, examines all compositions of the current $r_{i}$ into $|T|$ elements, with a runtime exponential in $r$. However, since $r$ is restricted to lower numbers, in practice, Algorithm S6 computes exact solutions efficiently.
Algorithm 5 Initialization of $s^{r}$ and $\mathcal{Q}$Input: $\bm{x},y,r,\hat{P},\mathcal{V}_{\mathcal{T}}$, Output: $s^{r}$ and $\mathcal{Q}$.
1:  $\mathcal{Q}\leftarrow\emptyset$triangleright initialize a priority queue
2:  $\mathcal{Q}\mathrm{.add}\big{(}(v_{1},\hat{P}(v_{1}\,|\,\bm{x}))\big{)}$triangleright tree root with corresponding probability mass
3:  while $\mathcal{Q}$ is not empty do
4:    $(v,\hat{p}_{v})\leftarrow\mathcal{Q}$.pop() $\triangleright$ pop the node with highest probability mass $\hat{P}(v\,|\,\bm{x})$
5:    $s^{r}(v)\leftarrow[\emptyset]\ast r$
6:    if $v$ is a leaf node then
7:       if $|\hat{\omega}(y;\bm{x})\cap v|\neq 0$ then
8:          if $\mathrm{pa}(v)\notin\mathcal{Q}$ then
9:             $\mathcal{Q}\mathrm{.add}(\mathrm{pa}(v))$
10:          end if
11:          $s^{r}(v)\leftarrow[v]\ast r$
12:       end if
13:    else
14:       for all children $v^{\prime}$ of $v$ do
15:          $\hat{p}_{v^{\prime}}\leftarrow\hat{p}_{v}\cdot\hat{P}(v^{\prime}\,|\,v,\bm{x})$triangleright compute probability mass of child node $v^{\prime}$ by means of (2)
16:          $\mathcal{Q}\mathrm{.add}\big{(}(v^{\prime},\hat{p}_{v^{\prime}})\big{)}$
17:       end for
18:    end if
19:  end while
20:  return $s^{r}$, $\mathcal{Q}$
Algorithm 6 Computation of $A_{r}(y;\bm{x})$ in (9) – Input: $\bm{x},y,r,\hat{P},\mathcal{V}_{\mathcal{T}}$
1:  $s^{r},\mathcal{Q}\leftarrow$ initialize $s^{r}$ and $\mathcal{Q}$ by means of Algorithm 5
2:  $\hat{Y}_{r}\leftarrow\emptyset$
3:  while $\mathcal{Q}$ is not empty do
4:    $v\leftarrow\mathcal{Q}$.pop()
5:    $T\leftarrow$ empty list
6:    for all children $v^{\prime}$ of $v$ do
7:       $T\leftarrow T+v^{\prime}$
8:    end for
9:    for $r_{i}=1$ to $r$ do
10:       if $|T|>r_{i}$ then
11:          $s^{r_{i}-1}(v)\leftarrow v$
12:       else
13:          $M\leftarrow$ all compositions of $r_{i}$ in $|T|$ elements
14:          $u^{*}\leftarrow+\infty$
15:          for $(i_{0},\ldots,i_{|T|-1})\in M$ do
16:             $s\leftarrow s^{i_{0}}(T[0])\cup\ldots\cup s^{i_{|T|-1}}(T[|T|-1])$
17:             if $|s|-\hat{p}_{s}<u^{*}$ then
18:               $s^{r_{i}-1}(v)\leftarrow s$
19:               $u^{*}\leftarrow|s|-\hat{p}_{s}$
20:             end if
21:          end for
22:       end if
23:    end for
24:    if $\mathrm{pa}(v)\notin\mathcal{Q}$ then
25:       $\mathcal{Q}\mathrm{.add}(\mathrm{pa}(v))$
26:    end if
27:  end while
28:  return $s^{r}(v_{1})$