### v3

# Conformal Prediction in Hierarchical Classification with Constrained Representation Complexity

Abstract
Conformal prediction has emerged as a widely used framework for constructing valid prediction sets in classification and regression tasks. In this work, we extend the split conformal prediction framework to hierarchical classification, where prediction sets are commonly restricted to internal nodes of a predefined hierarchy, and propose two computationally efficient inference algorithms. The first algorithm returns internal nodes as prediction sets, while the second one relaxes this restriction. Using the notion of representation complexity, the latter yields smaller set sizes at the cost of a more general and combinatorial inference problem. Empirical evaluations on several benchmark datasets demonstrate the effectiveness of the proposed algorithms in achieving nominal coverage.
1 Introduction
In multi-class classification, a classifier can be uncertain about the predicted class label for a given test instance. In such cases, it can be beneficial to return set-valued predictions, i.e. sets of classes rather than individual classes. This is particularly relevant in hierarchical classification, where the class space is organised in a hierarchical structure, such as in medical diagnosis, where diseases are organised in a tree structure based on the International Classification of Diseases (ICD) (World Health Organization,, 1978).
In hierarchical classification, set-valued predictions are often restricted to internal nodes of the hierarchy. Such predictions have a clear semantic interpretation and can be computed using efficient inference algorithms (Alex Freitas,, 2007; Bi and Kwok,, 2015; Rangwala and Naik,, 2017; Yang et al.,, 2017; Wang et al.,, 2021; Valmadre,, 2022). However, a restriction of this kind can affect efficiency when the classifier is uncertain between classes in different branches of the hierarchy. In these cases, predicting a single internal node typically yields large and uninformative sets.
To address this, some approaches allow any subset of classes as predictions, which improves flexibility but comes at the cost of higher semantic complexity and reduced interpretability (Oh,, 2017). More recently, a set-based utility maximisation framework has been proposed that makes a compromise between the two aforementioned extremes by introducing the notion of representation complexity (Mortier et al.,, 2022). The assumption behind this notion is that inner nodes of a hierarchy have semantic meaning and can therefore be used as meaningful stand-alone predictions. By limiting the number of internal nodes used to represent a prediction, the proposed framework allows users to control the trade-off between efficiency and interpretability.
Figure 1: A sample image of the Lotus corniculatus species from the PlantCLEF 2015 dataset (Göeau et al.,, 2015).
To illustrate the usefulness of representation complexity, consider the PlantCLEF 2015 dataset (Göeau et al.,, 2015). This dataset consists of over one hundred thousand images representing 1,000 species of trees, herbs, and ferns native to the Western European region and is characterised by significant class ambiguity, making accurate predictions on species level often impossible. In Figure 1, an example image is shown corresponding to the Lotus corniculatus species. Using our proposed algorithms, two set-valued predictions were computed, with a restricted representation complexity of $r=1$ and $r\leq 3$, respectively. The former yields the root node of the hierarchy, i.e. all 1,000 species, and is completely uninformative. Increasing the representation complexity to three improves efficiency and yields the prediction $\{\emph{Lotus corniculatus},\emph{Tulipa sylvestris},\emph{Ficaria verna}\}$, i.e., three visually-related species including the true label.
Contributions. In this work, we build further upon the concept of representation complexity and extend the split conformal prediction framework to the hierarchical multi-class classification setting. Conformal prediction provides valid and efficient set-valued predictions for any learning algorithm, without making any assumptions on the underlying data distribution (Vovk et al.,, 2005). In particular, given a trained (hierarchical) classifier, a desired coverage level $1-\alpha\in(0,1)$, calibration samples $\{(\bm{x}_{i},y_{i})\}_{i=1}^{N}\subset\mathcal{X}\times\mathcal{Y}$ and a test sample $(\bm{x}_{N+1},y_{N+1})$, drawn i.i.d. from an unknown distribution $P$, we would like to construct a prediction set $\hat{Y}\in 2^{\mathcal{Y}}$ with representation complexity $r$ for the test instance $(\bm{x}_{N+1},y_{N+1})$, such that the following marginal validity guarantee holds:
$\mathbb{P}\left[y_{N+1}\in\hat{Y}(\bm{x}_{N+1})\right]\geq 1-\alpha\,,\,\,\textrm{s.t.}\,\,R_{\mathcal{T}}(\hat{Y})\leq r\,$, (1)
[…]
After reviewing some essentials on hierarchical probabilistic classification and conformal prediction in Section 2, we propose in Section 3 two algorithms that construct valid prediction sets, in the sense of satisfying (1). The first algorithm is designed for the restrictive case $r=1$, while the second extends to $r>1$. Moreover, we show that our algorithms possess distribution-free finite sample guarantees. Finally, in Section 4, we evaluate the proposed algorithms in terms of coverage and efficiency on a range of benchmark datasets.
$v_{1}=\{1,2,3,4,5,6,7,8\}$$v_{2}=\{1,2,3,4\}$$v_{4}=\{1,2\}$ $\begin{subarray}{c}v_{8}=\{1\},\\ 0.15\end{subarray}$ $\begin{subarray}{c}v_{9}=\{2\},\\ 0.13\end{subarray}$ $v_{5}=\{3,4\}$ $\begin{subarray}{c}v_{10}=\{3\},\\ 0.08\end{subarray}$ $\begin{subarray}{c}v_{11}=\{4\},\\ 0.125\end{subarray}$ $v_{3}=\{5,6,7,8\}$$v_{6}=\{5,6\}$ $\begin{subarray}{c}v_{12}=\{5\},\\ 0.14\end{subarray}$ $\begin{subarray}{c}v_{13}=\{6\},\\ 0.125\end{subarray}$ $v_{7}=\{7,8\}$ $\begin{subarray}{c}v_{14}=\{7\},\\ 0.125\end{subarray}$ $\begin{subarray}{c}v_{15}=\{8\},\\ 0.125\end{subarray}$
Figure 2: An example of a tree structure $\mathcal{T}$ with class space $\mathcal{Y}=\{1,\ldots,8\}$ and nodes $\mathcal{V}_{\mathcal{T}}=\{v_{1},\ldots,v_{15}\}$. The root $v_{1}$ represents the class space $\mathcal{Y}$ and leaves $\{v_{8},\ldots,v_{15}\}$ represent the individual classes. The numbers in the leaf nodes represent the class probabilities for an instance $\bm{x}$.
2 Background, related work and formal description of existing concepts
2.1 Hierarchical multi-class classification
In a standard multi-class classification setting, one assumes that training and test data are i.i.d. according to an unknown distribution $P(\bm{x},y)$ on $\mathcal{X}\times\mathcal{Y}$, with $\mathcal{X}$ some instance space (e.g. images, documents, etc.) and $\mathcal{Y}=\{c_{1},\ldots,c_{K}\}$ a class space consisting of $K$ classes. Probabilistic multi-class classifiers estimate the conditional class probabilities $P(\cdot\,|\,\bm{x})$ over $\mathcal{Y}$, such that $0\leq P(c\,|\,\bm{x})\leq 1$ for all $c\in\mathcal{Y}$ and $\sum_{c\in\mathcal{Y}}P(c\,|\,\bm{x})=1\,.$ This distribution can be estimated using a wide range of well-known probabilistic models, such as logistic regression, linear discriminant analysis, gradient boosting trees or neural networks with a softmax output layer. At prediction time, set-valued prediction algorithms, such as split conformal prediction, return sets $\hat{Y}$ that are subsets of $\mathcal{Y}$. The probability mass of such a set can be computed as $P(\hat{Y}\,|\,\bm{x})=\sum_{c\in\hat{Y}}P(c\,|\,\bm{x})$.
In this paper, we will consider a hierarchical multi-class classification setting. Hence, we assume that a domain expert has defined a hierarchy over the class space, in the form of a tree structure $\mathcal{T}$ that in general contains $M$ nodes. $\mathcal{V}_{\mathcal{T}}=\{v_{1},\ldots,v_{M}\}$ will denote the set of nodes and every node identifies a set of classes. As special cases, the root $v_{1}$ represents the class space $\mathcal{Y}$, and the leaves represent individual classes; see Figure 2 for a simple example. In hierarchical classification, the strong restriction $\hat{Y}\in\mathcal{V}_{\mathcal{T}}$ is typically made for predicted sets; see e.g. Bi and Kwok, (2015). The probability mass of such a set can be computed as $P(v\,|\,\bm{x})=\sum_{c\in v}P(c\,|\,\bm{x})$, or by using the chain rule of probability:
$P(v\,|\,\bm{x})=\prod_{v^{\prime}\in\mathrm{Path}(v)}P(v^{\prime}\,|\,\mathrm{pa}(v^{\prime}),\bm{x})\,$, (2)
where $\mathrm{Path}(v)$ is a set of nodes on the path connecting the node $v$ and the root of the tree structure. $\mathrm{pa}(v)$ gives the parent of node $v$ and $P(v\,|\,\mathrm{pa}(v),\bm{x})$ represents the branch probability of node $v$ given its parent $\mathrm{pa}(v)$. Note that for the root node $v_{1}$ one has $P(v_{1}\,|\,\mathrm{pa}(v_{1}),\bm{x})=1$. In Figure 2, the branch probabilities of the root node $v_{1}$ are given by $P(v_{2}\,|\,v_{1},\bm{x})=0.485$ and $P(v_{3}\,|\,v_{1},\bm{x})=0.515$. In order to estimate the branch probabilities, one can train any multi-class probabilistic classifier in each internal node of the tree. Classical models of that kind include nested dichotomies (Fox,, 1997; Frank and Kramer,, 2004; Melnikov and Hüllermeier,, 2018), conditional probability estimation trees (Beygelzimer et al.,, 2009) and probabilistic classifier trees (Dembczyński et al.,, 2016). In neural networks with a hierarchical softmax output layer, all nodes are trained simultaneously (Morin and Bengio,, 2005).
2.2 Inference in hierarchical classification
[…]
$\mathcal{S}_{\mathcal{T}}(\hat{Y})=\left\{\hat{V}\subset\mathcal{V}_{\mathcal{T}}:\bigcup_{v_{i}\in\hat{V}}v_{i}=\hat{Y}\land\bigcap_{v_{i}\in\hat{V}}v_{i}=\emptyset\right\}\,.$
Then, we define the representation complexity of the prediction $\hat{Y}$ as
[…]
2.3 Randomized nested prediction sets with split conformal prediction
Now we will review a general procedure for valid set-valued predictions in flat classification (i.e. ignoring the hierarchy), following the work of Gupta et al., (2022); Romano et al., (2020); Angelopoulos et al., (2020). Compared to traditional conformal prediction—which starts from the notion of a nonconformity score—this procedure departs from a sequence of nested set-valued predictions, where the set size depends on a threshold $\tau$. Furthermore, an independent calibration set is used to tune the threshold $\tau$ such that $(1-\alpha)$-coverage is guaranteed on future test samples. The use of a single calibration set is better known as split conformal prediction and gives rise to computationally efficient valid set-valued predictions (Papadopoulos et al.,, 2002; Lei et al.,, 2018).
More formally, let $\{(\bm{x}_{i},y_{i})\}_{i=1}^{N}$ be an i.i.d. sequence of $N$ samples from the unknown distribution $P$, and assume that these samples were not used for model training. Let $\hat{Y}(\bm{x},u,\tau):\mathcal{X}\times[0,1]\times\mathbb{R}\rightarrow 2^{\mathcal{Y}}$ be a set-valued predictor. In the spirit of Romano et al., (2020); Angelopoulos et al., (2020), the second argument $u$ represents a random draw from a uniform distribution $\mathcal{U}(0,1)$ and is included to allow for randomized prediction sets. The third argument $\tau$ is a threshold that controls the size of the predicted set. Assume that the sets are indexed by $\tau$ and are nested:
$\hat{Y}(\bm{x},u,\tau_{1})\subseteq\hat{Y}(\bm{x},u,\tau_{2})\qquad\text{if}\quad\tau_{1}\leq\tau_{2}\,.$ (4)
In order to find the optimal threshold $\tau^{*}$ that guarantees $(1-\alpha)$-coverage, one needs to find the smallest $\tau$ such that the predicted set $\hat{Y}(\bm{x},u,\tau)$ contains at least $\lceil(N+1)(1-\alpha)\rceil$ samples:
$\tau^{*}=\inf\big\{\tau\in[0,1]:|\{i:y_{i}\in\hat{Y}(\bm{x}_{i},u_{i},\tau)\}|\\ \geq\lceil(1-\alpha)(N+1)\rceil\big\}\,.$ (5)
The set-valued predictor $\hat{Y}(\bm{x}_{N+1},u_{N+1},\tau^{*})$, with $\tau^{*}$ in (5), has marginal validity guarantees, as shown by the theorem below.
Theorem 2.1 (Marginal validity of nested conformal prediction (Angelopoulos et al.,, 2020)).
Assume an exchangeable sequence $\{(\bm{x}_{i},y_{i},u_{i})\}_{i=1}^{N+1}$ and let $\hat{Y}(\bm{x},u,\tau)$ be a set-valued predictor that satisfies (4). Furthermore, assume that $\exists\tau\in\mathbb{R}:\hat{Y}(\bm{x},u,\tau)=\mathcal{Y}$. Then, for $\tau^{*}$ in (5) and any $\alpha\in(0,1)$, the following marginal coverage guarantee holds:
$1-\alpha\leq\mathbb{P}\big[y_{N+1}\in\hat{Y}(\bm{x}_{n+1},u_{N+1},\tau^{*})\big]\,.$
In the literature, several conformal prediction methods can be found that satisfy (4). For example, Romano et al., (2020) propose the following set-valued predictor, called adaptive prediction sets (APS):
$\hat{Y}_{APS}(\bm{x},u,\tau)=\big\{y\in\mathcal{Y}:\hat{\rho}(y;\bm{x})+u\cdot\hat{P}(y\,|\,\bm{x})\leq\tau\big\}\,$, (6)
with $\hat{\rho}(y;\bm{x})=\sum_{y^{\prime}\in\mathcal{Y}}\hat{P}(y^{\prime}\,|\,\bm{x})\mathbbm{1}_{\hat{P}(y^{\prime}\,|\,\bm{x})>\hat{P}(y\,|\,\bm{x})}$ the probability mass of the labels more likely than $y$. At the heart of their method is the randomization term $u\cdot\hat{P}(y\,|\,\bm{x})$, which is used to achieve exact nominal coverage. In addition, by means of the cumulative distribution $\hat{\rho}(y;\bm{x})$, this method allows to adapt effectively to complex data distributions, achieving a better conditional coverage compared to alternative approaches that rely on the mode of the distribution, such as the least ambiguous classifier proposed by Sadinle et al., (2019). Angelopoulos et al., (2020) propose a variant of the above method, called the regularized adaptive prediction sets (RAPS) method. By introducing a regularization term in (6), small set sizes are encouraged, especially when facing noisy probability estimates for classes with low probability.
Alternative conformal prediction methods have been introduced, focusing on minimizing set size (Sadinle et al.,, 2019; Gao et al.,, 2025) and achieving approximate coverage across the entire feature space (Foygel Barber et al.,, 2021; Cauchois et al.,, 2021; Romano et al.,, 2019). Goren et al., (2024) and Angelopoulos et al., (2023) propose conformal prediction methods that control the risk with respect to bounded non-increasing loss functions. Interestingly, these methods are also applicable to the setting of hierarchical classification, e.g. through the use of the tree-distance loss, introduced by Bi and Kwok, (2015) as a way of evaluating set-valued predictions in multi-label classification. However, set-valued predictions of that kind lack meaningful and practical interpretation, compared to methods that rely on the traditional notion of coverage. Furthermore, these methods are not directly applicable to our work, as they do not allow the incorporation of representation complexity, given the properties of the loss functions above.
Gao et al., (2025) recently proposed a novel conformal predictor for regression, by implementing a dynamic programming algorithm that finds a union of $k$-intervals on the real line. When the output distribution is multi-modal, a union of $k$-intervals can be more informative and efficient compared to a single prediction interval (as obtained with most regression methods). Our work has a similar motivation, but adopted to hierarchical classification: when the classifier is uncertain between classes in different branches of the hierarchy, restricting the representation complexity to one is overly restrictive, while a higher complexity yields more efficient predictions.
3 Proposed methods
Building upon the concepts in Section 2.3, we are now ready to discuss two algorithms that provide valid set-valued predictions for hierarchical classification that satisfy (1) above. The first algorithm provides marginal validity guarantees in classical hierarchical classification settings, i.e. where sets are restricted to nodes of the tree structure, thus $R_{\mathcal{T}}(\hat{Y})=1$. The second algorithm provides marginal validity guarantees in restricted set-valued prediction settings, i.e. where the representation complexity of the predicted set is restricted to $R_{\mathcal{T}}(\hat{Y})\leq r$ for a user-defined value $r$. Note that traditional conformal prediction in flat classification is obtained for the second algorithm when there is no restriction on the representation complexity (e.g. $r=K$).
3.1 Conformal restricted set-valued prediction
Algorithm 1 CRSVP calibration – Input: $\{(\bm{x}_{i},y_{i},u_{i})\}_{i=1}^{N},\hat{P},\mathcal{V}_{\mathcal{T}}$, Output: Threshold in (5).
1:for $i=1,\ldots,N$ do
2:  $\hat{Y}\leftarrow\operatorname*{\arg\max}_{c\in\mathcal{Y}}\hat{P}(c\,|\,\bm{x}_{i})$
3:  $\hat{p}_{\hat{Y}}\leftarrow\hat{P}(\hat{Y}\,|\,\bm{x}_{i}),\hat{p}_{\hat{Y}^{\prime}}\leftarrow 0$
4:  while $y_{i}\notin\hat{Y}$ do
5:   $\hat{p}_{\hat{Y}^{\prime}}\leftarrow\hat{p}_{\hat{Y}}$, $\hat{p}_{\hat{Y}}\leftarrow\hat{P}(\mathrm{pa}(\hat{Y})\,|\,\bm{x}_{i})$
[…]
7:  end while
8:  $\tau_{i}\leftarrow\hat{p}_{\hat{Y}}-u_{i}\cdot(\hat{p}_{\hat{Y}}-\hat{p}_{\hat{Y}^{\prime}})$
9:end for
[…]
11:return $\tau^{*}$
Algorithm 2 CRSVP inference – Input: $\bm{x},\tau^{*},u,\hat{P},\mathcal{V}_{\mathcal{T}}$, Output: Set-valued prediction in (7).
[…]
3:while $\mathrm{pa}(\hat{Y}^{\prime})\neq\emptyset$ do
4:  if $\hat{p}_{\hat{Y}}\geq\tau^{*}$ then
5:   break
6:  end if
7:  $\hat{p}_{\hat{Y}^{\prime}}\leftarrow\hat{p}_{\hat{Y}}$, $\hat{p}_{\hat{Y}}\leftarrow\hat{P}(\mathrm{pa}(\hat{Y})\,|\,\bm{x})$
[…]
9:end while
10:if $\hat{p}_{\hat{Y}}-u\cdot(\hat{p}_{\hat{Y}}-\hat{p}_{\hat{Y}^{\prime}})\geq\tau^{*}$ then
11:  return $\hat{Y}^{\prime}$
12:else
13:  return $\hat{Y}$
14:end if
The first approach, called conformal restricted set-valued prediction (CRSVP), predicts sets that are restricted to nodes of the hierarchy, thus having representation complexity $R_{\mathcal{T}}(\hat{Y})=1$. Assume one has fitted a classifier using a training dataset as described in Section 2.1, and let $\hat{y}(\bm{x})$ denote the mode of the distribution $\hat{P}(.\,|\,\bm{x})$, i.e. the leaf node with highest probability mass. For our first approach, we consider the following set-valued predictor:
$\hat{Y}_{1}(\bm{x},u,\tau)=\operatorname*{\arg\max}_{\hat{Y}\in\mathrm{Path}(\hat{y}(\bm{x}))}\big\{|\hat{Y}|:\hat{P}(\hat{Y}\,|\,\bm{x})+\\ u\cdot\hat{P}(\mathrm{pa}(\hat{Y})\setminus\hat{Y}\,|\,\bm{x})\leq\tau\big\}\,$, (7)
with $\mathrm{Path}(v)$ the path to the root, and $\mathrm{pa}(v)$ the parent of a node $v$, as defined in Section 2. $\hat{P}(\hat{Y}\,|\,\bm{x})$ can be computed using (2), but also observe that the probability mass of an internal node $v$ corresponds to $\hat{P}(v\,|\,\bm{x})=\sum_{c\in v}\hat{P}(c\,|\,\bm{x})$, i.e. the sum of the probability masses of the leaf nodes that are descendants of $v$. However, computing the probability mass of an internal node and its associated set of classes in this way is computationally less efficient than using the chain rule in (2), when hierarchical classifiers have been fitted during training.
It is clear that the set-valued predictor in (7) satisfies the nestedness property in (4). Indeed, starting from the mode of the distribution $\hat{y}(\bm{x})$, every node on the path connecting that node and the root is nested by definition of the hierarchical tree structure $\mathcal{T}$. In line with Angelopoulos et al., (2020), the first term increases as we move up the tree, while the second term contains a random draw from the uniform distribution $\mathcal{U}(0,1)$, for handling discrete jumps in probability mass when following the path towards the root. The randomization is needed to prevent over-coverage.
The algorithm for calculating the threshold in (5) using the calibration dataset is presented in Algorithm 1. This algorithm first computes the mode of the estimated class probabilities for each instance in the calibration dataset. For small hierarchies, this can be done using exhaustive search, but more efficient inference algorithms have been developed when the computational cost becomes a burden (Dembczyński et al.,, 2012; Kumar et al.,, 2013; Mena Waldo et al.,, 2015). Subsequently, starting from the mode, Algorithm 1 computes for every calibration instance the internal node, on the path to the root, that also includes the true class label. Nonconformity scores that incorporate a randomization component are constructed for these internal nodes, and the final threshold is deduced from the scores. Algorithm 1 has a worst case $\mathcal{O}(N\,K)$ time complexity when exhaustive search is applied in the first step.
The pseudocode for computing the set-valued prediction for a new test instance, as defined in (7), is shown in Algorithm 2. Given a hierarchical classifier, the computational complexity during test time is given by $\mathcal{O}(\log K)$.
3.2 Conformal set-valued prediction with representation complexity
Algorithm 3 CRSVP-$r$ calibration – Input: $\{(\bm{x}_{i},y_{i},u_{i})\}_{i=1}^{N},r,\hat{P},\mathcal{V}_{\mathcal{T}}$, Output: Threshold in (5).
1:for $i=1,\ldots,N$ do
2:  $k\leftarrow 1$
3:  $\hat{Y}^{(k-1)}\leftarrow\emptyset,\hat{Y}^{(k)}\leftarrow\operatorname*{\arg\max}_{c\in\mathcal{Y}}\hat{P}(c\,|\,\bm{x}_{i})$
4:  $\hat{p}_{\hat{Y}^{(k-1)}}\leftarrow 0,\hat{p}_{\hat{Y}^{(k)}}\leftarrow\hat{P}(\hat{Y}^{(k)}\,|\,\bm{x}_{i})$
5:  while $k<K$ do
6:   if $y_{i}\notin\hat{Y}^{(k)}$ then
7:    $\hat{Y}^{(k)^{\prime}}\leftarrow A_{r}(S_{k};\bm{x}_{i})$ (by means of Algorithm 5)
8:    $\hat{p}_{\hat{Y}^{(k)^{\prime}}}\leftarrow\hat{P}(\hat{Y}^{(k)^{\prime}}\,|\,\bm{x}_{i})$
9:    if $|\hat{Y}^{(k)^{\prime}}|\neq|\hat{Y}^{(k-1)}|$ then
10:     $\hat{Y}^{(k)},\hat{p}_{\hat{Y}^{(k)}}\leftarrow\hat{Y}^{(k)^{\prime}},\hat{p}_{\hat{Y}^{(k)^{\prime}}}$
11:    end if
12:    $k\leftarrow k+1$
13:   else
14:    break
15:   end if
16:  end while
17:  $\tau_{i}\leftarrow\hat{p}_{\hat{Y}^{(k)}}-u_{i}\cdot(\hat{p}_{\hat{Y}^{(k)}}-\hat{p}_{\hat{Y}^{(k-1)}})$
18:end for
[…]
20:return $\tau^{*}$
Algorithm 4 CRSVP-$r$ inference – Input: $\bm{x},\tau^{*},u,r,\hat{P},\mathcal{V}_{\mathcal{T}}$, Output: Set-valued prediction in (9).
$1k\leftarrow$ 1
[…]
4:while $k<K$ do
5:  if $y^{(k)}\notin\hat{Y}^{(k-1)}$ then
6:   $\hat{Y}^{(k)^{\prime}}\leftarrow A_{r}(S_{k};\bm{x})$ (by means of Algorithm 5)
7:   $\hat{p}_{\hat{Y}^{(k)^{\prime}}}\leftarrow\hat{P}(\hat{Y}^{(k)^{\prime}}\,|\,\bm{x})$
8:   if $|\hat{Y}^{(k)^{\prime}}|\neq|\hat{Y}^{(k-1)}|$ then
9:    $\hat{Y}^{(k)},\hat{p}_{\hat{Y}^{(k)}}\leftarrow\hat{Y}^{(k)^{\prime}},\hat{p}_{\hat{Y}^{(k)^{\prime}}}$
10:   end if
11:   if $\hat{p}_{\hat{Y}^{(k)}}\geq\tau^{*}$ then
12:    break
13:   end if
14:   $k\leftarrow k+1$
15:  end if
16:end while
17:if $\hat{p}_{\hat{Y}^{(k)}}-u\cdot(\hat{p}_{\hat{Y}^{(k)}}-\hat{p}_{\hat{Y}^{(k-1)}})\geq\tau^{*}$ then
18:  return $\hat{Y}^{(k-1)}$
19:else
20:  return $\hat{Y}^{(k)}$
21:end if
Algorithm 5 Dynamic programming solution for set of lowest common ancestors – Input: $\bm{x},S_{k},r,\hat{P},\mathcal{V}_{\mathcal{T}},A_{r}(S_{k-1};\bm{x})$, Output: $A_{r}(S_{k};\bm{x})$
[…]
2:for all $v\in\mathcal{V}_{\mathcal{T}}$ do
3:  $s^{1}(v)\leftarrow\emptyset,\ldots,s^{r}(v)\leftarrow\emptyset$
4:  if $|v|=1\;\wedge\;v\cap S_{k}\neq\emptyset$ then
5:   if $\mathrm{pa}(v)\notin\mathcal{Q}$ then
6:    $\mathcal{Q}\mathrm{.add}(\mathrm{pa}(v))$
7:   end if
8:   $s^{1}(v)\leftarrow v,\ldots,s^{r}(v)\leftarrow v$ $\triangleright$ initialize leaves that overlap with $S_{k}$
9:  end if
10:end for
11:while $\mathcal{Q}$ is not empty do
12:  $v\leftarrow\mathcal{Q}$.pop()
13:  $T\leftarrow\{v^{\prime}:v^{\prime}\in\mathrm{ch}(v)\wedge v^{\prime}\cap S_{k}\neq\emptyset\}$ $\triangleright$ visit all children that overlap with $S_{k}$
14:  for $r_{i}=1$ to $r$ do
15:   if $|T|>r_{i}$ then
16:    $s^{r_{i}}(v)\leftarrow v$
17:   else
18:    $M\leftarrow$ all compositions of $r_{i}$ into $|T|$ elements
19:    $u^{*}\leftarrow+\infty$
20:    for $(i_{0},\ldots,i_{|T|-1})\in M$ do
21:     $s\leftarrow s^{i_{0}}(T[0])\cup\cdots\cup s^{i_{|T|-1}}(T[|T|-1])$
22:     if $|s|-\hat{p}_{s}<u^{*}$ then
23:      if $v$ is the root then
24:       if $A_{r}(S_{k-1};\bm{x})\subseteq s$ then
25:        $s^{r_{i}}(v)\leftarrow s$
26:        $u^{*}\leftarrow|s|-\hat{p}_{s}$
27:       end if
28:      else
29:       $s^{r_{i}}(v)\leftarrow s$
30:       $u^{*}\leftarrow|s|-\hat{p}_{s}$
31:      end if
32:     end if
33:    end for
34:   end if
35:  end for
36:  if $\mathrm{pa}(v)\notin\mathcal{Q}\wedge\mathrm{pa}(v)\neq\emptyset$ then
37:   $\mathcal{Q}\mathrm{.add}(\mathrm{pa}(v))$
38:  end if
39:end while
40:return $s^{r}(v_{1})$
It should be obvious that sets of representation complexity one quickly become very big, since they correspond to internal nodes of the hierarchy that exhibit a predefined coverage level. Therefore, we present a second approach, called conformal set-valued prediction with representation complexity (CRSVP-$r$), that relaxes the representation complexity constraint to $R_{\mathcal{T}}(\hat{Y})\leq r$ with $r$ a user-defined parameter. Let $S_{k}\equiv\{y^{(1)},\ldots,y^{(k)}\}$ denote the top-$k$ classes for $\bm{x}$ when sorting the classes according to $\hat{P}(y\,|\,\bm{x})$, similar as in (6). We define the following sequence of optimization problems for any $k\in\{1,...,K\}$:
$A_{r}(S_{k};\bm{x})=\operatorname*{\arg\min}_{\begin{subarray}{c}\hat{Y}\in 2^{\mathcal{Y}}:R_{\mathcal{T}}(\hat{Y})\leq r,\\ A_{r}(S_{k-1};\bm{x})\,\cup\,y^{(k)}\subseteq\hat{Y}\end{subarray}}|\hat{Y}|-\hat{P}(\hat{Y}\,|\,\bm{x})\,$, (8)
with $A_{r}(S_{1};\bm{x})=y^{(1)}$. Remark that these optimization problems have to be solved in the correct order, because of the nestedness condition. For fixed $k$ we refer to this optimization problem as finding the set of common ancestors. It can be thought of as a variant of the well-known lowest common ancestor problem (Aho et al.,, 1973; Harel and Tarjan,, 1984; Bender and Farach-Colton,, 2000; Dash et al.,, 2013), i.e. instead of considering a single common ancestor of a set of leaves in a tree, we are looking for a set of $r$ non-overlapping ancestors whose descendants form a set of representation complexity $r$. Moreover, we introduce the nested set condition, and the second term in the objective function is also not considered in the lowest common ancestor problem. Those two changes are needed to make sure that (1) the solutions for increasing $k$ are nested, and (2) the solution for fixed $k$ is unique. Since probabilities are bounded between zero and one, the second term only plays a role for sets that have the same cardinality (i.e. the two terms are lexicographically ordered). As an example, consider the hierarchy depicted in Figure 2. For $S_{3}=\{1,2,5\}$, the lowest common ancestor is given by $A_{1}(S_{3};\bm{x})=\{1,2,3,4,5,6,7,8\}$. The lowest set of common ancestors with representation complexity two is $A_{2}(S_{3};\bm{x})=\{1,2,5\}$, illustrating that set size can be drastically reduced by increasing the representation complexity.
Let $\hat{Y}^{(1)},\ldots,\hat{Y}^{(K)}$ denote the ordered sequence of unique lowest common ancestors $A_{r}(S_{1};\bm{x}),\ldots,A_{r}(S_{K};\bm{x})$. We use this sequence to introduce the following set-valued predictor:
$\hat{Y}_{\leq r}(\bm{x},u,\tau)=\operatorname*{\arg\max}_{\hat{Y}^{(k)}\in\{\hat{Y}^{(1)},\ldots,\hat{Y}^{(K)}\}}\big\{|\hat{Y}^{(k)}|:\\ \hat{P}(\hat{Y}^{(k-1)}\,|\,\bm{x})+u\cdot\hat{P}(\hat{Y}^{(k)}\setminus\hat{Y}^{(k-1)}\,|\,\bm{x})\leq\tau\big\}\,.$ (9)
Due to the nested set constraint in (8), the set-valued predictor in (9) satisfies the nestedness property in (4). Thus, this set-valued predictor automatically has the classical marginal coverage guarantees of conformal prediction. Similar as for the first approach, the randomization term $u\cdot\hat{P}(\hat{Y}^{(k)}\setminus\hat{Y}^{(k-1)}\,|\,\bm{x})$ is again included to prevent over-coverage.
Let us now discuss the algorithmic aspects of computing the set of lowest common ancestors in (8). At first impression, the combinatorial optimization problem in (8) looks challenging, but it can be solved in an efficient manner with a divide-and-conquer algorithm that is described in Algorithm 5.
For every internal node of the hierarchy, the optimization problem can be broken down into simpler subproblems of the same type. As an example, consider in a predefined hierarchy an internal node $v^{\prime}$ that has four children $(v^{\prime}_{1},v^{\prime}_{2},v^{\prime}_{3},v^{\prime}_{4})$ containing at least one element of $S_{k}$, and assume that we aim to find the set of lowest common ancestors with representation complexity three for a set of classes $S_{k}$ that are all descendants of $v^{\prime}$. As potential divisions of the representation complexity, one could find one common ancestor as descendants of children $v^{\prime}_{1}$, $v^{\prime}_{3}$, and $v^{\prime}_{4}$, or as descendants of children $v^{\prime}_{1}$, $v^{\prime}_{2}$, and $v^{\prime}_{4}$, or as two descendants of $v^{\prime}_{1}$ combined with one descendant of $v^{\prime}_{4}$, etc. In fact, many combinations are possible. Speaking in formal combinatorial terminology, one has to consider for the node $v$ all compositions of the integer three into four elements. For each of these compositions, the optimization problem can be recursively divided into smaller problems with lower representation complexity, where the node $v$ is replaced by each of its non-zero children in the composition. However, implementing such a strategy in a recursive manner would heavily blow up the computations, since each of the smaller subproblems would need to be solved many times.
A dynamic programming implementation, which avoids recursion by solving the smaller subproblems in a bottom-up manner, before tackling the more challenging optimization problems higher up in the hierarchy, is able to solve (8) in an efficient manner for small values of $r$. The pseudocode of such an implementation is described in Algorithm 5. For every internal node, $r$ local optimization problems are solved by varying the local representation complexity $r_{i}$ from one to $r$, and the solutions of these optimization problems are stored. By visiting children before their parents get analysed, one can guarantee that all needed quantities are known when the different compositions in an internal node need to be investigated. The most critical step in Algorithm 5 is line 20, where all compositions of the current representation complexity $r_{i}$ into $|T|$ elements are considered (with $T$ the set of all children that contain at least one element of $S_{k}$). Strictly speaking, this step has a runtime that is exponential in $r$. However, in practical situations, one would only be interested in sets with a representation complexity lower than five, so Algorithm 5 in general computes the exact solution to (8) in an efficient manner.
The pseudocode for calculating the threshold in (5) and set-valued predictions in (9) is presented in Algorithm 3 and 4. The computational complexity during test time is dominated by the computation of the set of lowest common ancestors in (8), which needs to be computed for every $k$ until the stopping criterion in Algorithm 4 is satisfied. Furthermore, the worst-case time complexity of calculating the set of lowest common ancestors in Algorithm 5 is upper bounded by $\mathcal{O}(K^{2}r^{d})$, with $d$ the maximum out-degree of the tree $\mathcal{T}$. This is small for regimes that are practically useful (fixed $r\leq 3$ and moderate $d$).
4 Experimental results
Table 1: Summary statistics and top-1 performance for all datasets. Notation: $K$ – number of classes, NN – number of samples, Top-1 acc. – top-1 accuracy of classifier.
Dataset $K$ $N_{train}$ $N_{cal}$ $N_{test}$ Top-1 acc.
CIFAR-10 10 50000 5000 5000 0.8817
AMB 93 12781 1918 1917 0.8503
Caltech-101 97 4338 2169 2169 0.9039
DBpedia 219 276945 30397 30397 0.9388
Caltech-256 256 14890 7445 7445 0.7578
PlantCLEF 2015 1000 91758 10723 10723 0.4156
In this section, different set-valued predictors are compared for solving problem (1) across six benchmark datasets, focusing on coverage, efficiency, and representation complexity. Summary statistics for these datasets are presented in Table 1. For all datasets, a predefined hierarchy $\mathcal{T}$ is extracted from the provided hierarchical labels. For detailed information about the experimental setup, we refer to Section A in the technical appendix.
Table 2: Results for CIFAR-10, AMB, Caltech-101, DBpedia, Caltech-256 and PlantCLEF 2015. Coverage, efficiency, and representation complexity for the following unrestricted set-valued predictors: LAC, NPS, APS, and restricted set-valued predictors: NCRSVP, CRSVP, NCRSVP-33, and CRSVP-$3$. The confidence level is set to 90%, and calibration and test sets are resampled 10 times.
Alg. LAC NPS APS NCRSVP CRSVP NCRSVP-3 CRSVP-3
Dataset
CIFAR-10 Cov. $0.899\pm 0.005$ $0.997\pm 0.001$ $0.899\pm 0.003$ $1.000\pm 0.000$ $0.899\pm 0.005$ $0.997\pm 0.001$ $0.899\pm 0.003$
Size $1.473\pm 0.023$ $5.125\pm 0.058$ $1.849\pm 0.019$ $10.00\pm 0.000$ $3.899\pm 0.049$ $5.861\pm 0.064$ $1.946\pm 0.025$
R.C. $1.451\pm 0.021$ $3.552\pm 0.017$ $1.824\pm 0.015$ $1.000\pm 0.000$ $1.000\pm 0.000$ $2.368\pm 0.008$ $1.691\pm 0.009$
AMB Cov. $0.899\pm 0.010$ $1.000\pm 0.000$ $0.900\pm 0.009$ $1.000\pm 0.000$ $0.900\pm 0.009$ $1.000\pm 0.000$ $0.899\pm 0.009$
Size $1.128\pm 0.020$ $23.98\pm 1.195$ $1.685\pm 0.055$ $50.85\pm 1.304$ $4.856\pm 0.261$ $39.26\pm 1.574$ $2.184\pm 0.063$
R.C. $1.132\pm 0.019$ $17.38\pm 0.702$ $1.776\pm 0.051$ $1.000\pm 0.000$ $1.000\pm 0.000$ $1.829\pm 0.009$ $1.394\pm 0.018$
Caltech-101 Cov. $0.900\pm 0.007$ $1.000\pm 0.000$ $0.900\pm 0.006$ $1.000\pm 0.000$ $0.901\pm 0.005$ $1.000\pm 0.000$ $0.901\pm 0.006$
Size $0.920\pm 0.008$ $96.54\pm 0.065$ $1.165\pm 0.015$ $96.25\pm 0.049$ $4.400\pm 0.254$ $63.61\pm 0.517$ $1.784\pm 0.086$
[…]
DBpedia Cov. $0.899\pm 0.001$ $1.000\pm 0.000$ $0.901\pm 0.002$ $0.999\pm 0.000$ $0.901\pm 0.002$ $0.999\pm 0.000$ $0.901\pm 0.001$
Size $0.931\pm 0.002$ $59.17\pm 0.272$ $11.33\pm 0.287$ $162.9\pm 0.492$ $26.62\pm 0.417$ $127.9\pm 0.501$ $17.71\pm 0.423$
R.C. $1.000\pm 0.000$ $53.18\pm 0.257$ $11.90\pm 0.277$ $1.000\pm 0.000$ $1.000\pm 0.000$ $1.985\pm 0.004$ $1.380\pm 0.005$
Caltech-256 Cov. $0.900\pm 0.004$ $0.999\pm 0.000$ $0.901\pm 0.004$ $0.999\pm 0.000$ $0.900\pm 0.003$ $1.000\pm 0.000$ $0.901\pm 0.003$
Size $1.931\pm 0.040$ $62.25\pm 1.330$ $3.640\pm 0.099$ $208.4\pm 1.780$ $44.69\pm 1.252$ $163.6\pm 2.070$ $20.30\pm 0.830$
R.C. $1.926\pm 0.040$ $50.12\pm 0.996$ $3.680\pm 0.098$ $1.000\pm 0.000$ $1.000\pm 0.000$ $1.632\pm 0.005$ $1.498\pm 0.009$
PlantCLEF 2015 Cov. $0.899\pm 0.003$ $0.970\pm 0.001$ $0.899\pm 0.003$ $1.000\pm 0.000$ $0.900\pm 0.002$ $1.000\pm 0.000$ $0.899\pm 0.004$
Size $25.50\pm 0.450$ $123.6\pm 2.026$ $44.12\pm 0.994$ $998.9\pm 0.224$ $520.9\pm 4.745$ $995.3\pm 0.492$ $389.7\pm 5.898$
R.C. $24.33\pm 0.419$ $104.4\pm 1.560$ $40.19\pm 0.836$ $1.000\pm 0.000$ $1.000\pm 0.000$ $1.006\pm 0.001$ $1.632\pm 0.010$
In particular, we will compare various set-valued predictors that solve problem (1) across six datasets listed in Table 1: CIFAR-10 (Krizhevsky et al.,, 2010), Caltech-101 (Li et al.,, 2003), Caltech-256 (Griffin et al.,, 2007), PlantCLEF 2015 (Göeau et al.,, 2015), the Allen Mouse Brain (AMB) single-cell transcriptomics dataset (Tasic et al.,, 2018) and DBPedia (Ofer,, 2019). The first four datasets are image datasets from the computer vision domain. The AMB dataset contains gene expression profiles from the mouse neocortex along with corresponding cell types and a hierarchical structure. The DBpedia text dataset contains extracted structured textual content from Wikipedia. Specifically, we evaluate the following set-valued predictors: CRSVP and CRSVP-$3$. Additionally, we demonstrate the usefulness of randomized prediction sets by considering the following naïve (N) set-valued predictors: NCRSVP and NCRSVP-$3$. These correspond to setting $u$ to zero when calculating the threshold in (5) and constructing the prediction sets in (7) and (9), respectively. Finally, we include results for three baseline methods that produce valid set-valued predictions in flat classification (i.e. ignoring the hierarchy): the least ambiguous classifier (LAC), as proposed by Sadinle et al., (2019); adaptive prediction sets (APS), as defined in (6); and naïve prediction sets (NPS), which correspond to APS without randomization (i.e. setting $u$ to zero). The results presented below are obtained using a flat classifier. We do not use hierarchical classifiers, as our dynamic programming approach relies on a bottom-up traversal of the tree, and incorporating hierarchical classifiers would increase computational complexity. Moreover, Mortier et al., (2022) report no significant improvement in predictive performance when using hierarchical classifiers compared to flat ones. Nevertheless, hierarchical classifiers can still be applied within our framework, as discussed in Section 2.1. During inference, we resample the calibration and test sets 10 times and use a confidence level of 90%.
The most important results are summarized in Table 2. For each experiment, we report (marginal) coverage, efficiency, and representation complexity. Coverage is defined as the proportion of samples for which the true class is included in the prediction set. Efficiency is defined as the average size of the set-valued prediction. The results clearly indicate that naïve set-valued predictors fail to deliver prediction sets with exact coverage, thereby highlighting the importance of randomized prediction sets. Moreover, increasing the representation complexity generally improves efficiency, demonstrating its practical value. In extreme cases, when representation complexity is unrestricted, such as with LAC, NPS, and APS, optimal performance in terms of efficiency is observed. However, this comes at the cost of significantly increased representation complexity in the prediction sets, in particular for large $K$, which may not be practical when predictions need to adhere to a predefined hierarchy.
Note that imprecise predictions for low representation complexity are typically due to uncertainty in the predicted probabilities and/or characteristics of the hierarchical structure, such as inconsistency and depth. Specifically, when uncertainty spans distinct branches of the hierarchy, predictions tend to correspond to higher-level nodes for lower representation complexity. This suggests that the hierarchy may not be well-suited for certain samples. Additionally, the depth of the hierarchical structure also plays a critical role in hierarchical classification tasks with a large number of classes. In such cases, a shallow hierarchy can lead to imprecise predictions for low representation complexity, as internal nodes in shallow trees often have many children. For example, the PlantCLEF 2015 dataset, characterized by 1,000 classes and a shallow hierarchy (i.e. taxonomy consisting of a family, genus and species level), requires a higher representation complexity in order to improve efficiency.
Finally, higher representation complexity becomes valuable for datasets with many classes and high uncertainty. In this case, traditional hierarchical classification tends to return imprecise predictions, i.e. nodes higher up in the hierarchy. By restricting representation complexity, predictions can span multiple nodes lower in the hierarchy, thereby improving efficiency. This is demonstrated by comparing CRSVP and CRSVP-3 on the PlantCLEF 2015 dataset in Table 1 and Figure 3. Furthermore, Figure 3 reveals a clear trade-off between representation complexity and efficiency for the PlantCLEF 2015 dataset. In Section B of the technical appendix, we present additional results obtained for (N)CRSVP-$1$ and (N)CRSVP-$2$, and provide additional insights regarding conditional coverage and the relationship between representation complexity and efficiency.
Figure 3: Trade-off between representation complexity (log scale) and efficiency for PlantCLEF 2015. The confidence level is set to 90%, and calibration and test sets are resampled 10 times.
5 Conclusion
In this work, we extended the split conformal prediction framework to hierarchical classification by introducing two novel set-valued prediction algorithms. The first algorithm generates valid set-valued predictions restricted to single nodes within a predefined hierarchy. We argued that this restriction can be limiting in certain applications. To address this limitation, we introduced a second algorithm that relaxes this constraint by incorporating the notion of representation complexity. Empirical evaluations on multiple benchmark datasets demonstrate the effectiveness of the proposed algorithms in achieving exact nominal coverage.
The main purpose of bounding representation complexity is to ensure the interpretability and semantic meaningfulness of predictions. However, it may also have other benefits. In particular, we conjecture that it may serve the purpose of regularization and hence increase accuracy in cases where class probabilities are poorly estimated. In such cases, flat prediction sets tend to be scattered across the entire hierarchy, and this scattering is suppressed by bounding representation complexity. We plan to elaborate on this aspect more closely in future work. Another interesting direction for future research is to generalize our methods to more complex structures, such as directed acyclic graphs.
Acknowledgements
Yusuf Sale is supported by the DAAD program Konrad Zuse Schools of Excellence in Artificial Intelligence, sponsored by the Federal Ministry of Education and Research. Alireza Javanmardi gratefully acknowledges funding by the Klaus Tschira Stiftung (project 00.019.2024). Willem Waegeman received funding from the Flemish government under the Flanders AI research (FlAIR) program.
Supplementary Materials
Appendix A Experimental setup
We use a MobileNetV2 convolutional neural network (Sandler et al.,, 2018) pretrained on ImageNet (Krizhevsky et al.,, 2017), in order to obtain hidden representations for all image datasets. The cross-entropy loss is minimized using the Adam optimizer, with a learning rate of $1\times 10^{-5}$ and momentum set to 0.99. We set the number of epochs to 2 and 20, for the Caltech and other datasets, respectively. We train all models end-to-end on a GPU, by using the PyTorch library (Paszke et al.,, 2019) and infrastructure with the following specifications:
  • CPU: Intel i7-6800K 3.4 GHz (3.8 GHz Turbo Boost)
  • GPU: NVIDIA GTX 1080 Ti 11GB
  • RAM: 64GB DDR4-2666
For the AMB and DBpedia dataset, we use the same model and training procedure as described in Theunissen et al., (2024) and Mortier et al., (2021), respectively.
Appendix B Additional results
In this section, we include additional results to those presented in Table 2, namely, results for (N)CRSVP-$1$ and (N)CRSVP-$2$. In addition, we give insights into potential conditional coverage violations by examining the Pearson correlation between coverage and prediction set size. A strong correlation may indicate a potential violation of conditional coverage (Feldman et al.,, 2021). While this metric serves as a useful proxy for evaluating conditional coverage, it is not definitive, as a correlation of zero does not necessarily guarantee conditional coverage, as shown in Rossellini et al., (2024). The results are summarized in Table C3, C4 and C5.
First, note that CRSVP is different from CRSVP-$1$. The restriction is the same, however, the set of solutions for both algorithms is different: For CRSVP, starting from the mode of the distribution, the solution space is given by the mode and its ancestors. For CRSVP-$1$, the solution space is given by line 7 and 6 in Algorithm 2 and 3, respectively, namely, the most common ancestor (8) for the top-1, top-2, top-3, etc., classes. In most cases, we observe low correlation for (N)CRSVP and (N)CRSVP-$r$, reflecting strong performance regarding conditional coverage.
Finally, in Figure C4, we visually explore the relationship between representation complexity and set size for the unrestricted set-valued predictors (LAC, APS) and restricted set-valued predictors (CRSVP, CRSVP-$r$). As is clear from both Table C3, C4, C5 and Figure C4, increasing the representation complexity results in better efficiency.
Table C3: Additional results for CIFAR-10 and AMB. Coverage, efficiency, representation complexity and Pearson correlation values for the following unrestricted set-valued predictors: LAC, NPS, APS, and restricted set-valued predictors: NCRSVP, CRSVP, NCRSVP-$r$, and CRSVP-$r$. The confidence level is set to 90%, and calibration and test sets are resampled 10 times.
Dataset CIFAR-10 AMB
Alg. Cov. Size Repr. comp. Corr. Cov. Size Repr. comp. Corr.
LAC $0.899\pm 0.005$ $1.473\pm 0.023$ $1.451\pm 0.021$ $0.07\pm 0.011$ $0.899\pm 0.010$ $128\pm 0.020 1.1$$132\pm 0.019$ $0.060\pm 0.025$
NPS $0.997\pm 0.001$ $5.125\pm 0.058$ $3.552\pm 0.017$ $0.016\pm 0.010$ $1.000\pm 0.000$ $23.98\pm 1.195$ $17.38\pm 0.702$ -
APS $0.899\pm 0.003$ $1.849\pm 0.019$ $1.824\pm 0.015$ $0.111\pm 0.014$ $0.900\pm 0.009$ $1.685\pm 0.055$ $1.776\pm 0.051$ $0.177\pm 0.015$
NCRSVP $1.000\pm 0.000$ $10.00\pm 0.000$ $1.000\pm 0.000$ - $1.000\pm 0.000$ $50.85\pm 1.304$ $1.000\pm 0.000$ -
CRSVP $0.899\pm 0.005$ $3.899\pm 0.049$ $1.000\pm 0.000$ $0.134\pm 0.009$ $0.900\pm 0.009$ $4.856\pm 0.261$ $1.000\pm 0.000$ $0.137\pm 0.008$
NCRSVP-1 $1.000\pm 0.000$ $10.00\pm 0.000$ $1.000\pm 0.000$ - $1.000\pm 0.000$ $51.13\pm 1.338$ $1.000\pm 0.000$ -
CRSVP-1 $0.898\pm 0.005$ $3.580\pm 0.050$ $1.000\pm 0.000$ $0.207\pm 0.005$ $0.900\pm 0.008$ $4.890\pm 0.266$ $1.000\pm 0.000$ $0.148\pm 0.007$
NCRSVP-2 $0.998\pm 0.001$ $6.775\pm 0.047$ $1.616\pm 0.004$ $0.028\pm 0.009$ $1.000\pm 0.000$ $43.35\pm 1.578$ $1.511\pm 0.011$ -
CRSVP-2 $0.899\pm 0.005$ $2.279\pm 0.036$ $1.457\pm 0.005$ $0.105\pm 0.012$ $0.900\pm 0.009$ $2.666\pm 0.117$ $1.248\pm 0.011$ $0.124\pm 0.007$
NCRSVP-3 $0.997\pm 0.001$ $5.861\pm 0.064$ $2.368\pm 0.008$ $0.024\pm 0.009$ $1.000\pm 0.000$ $39.26\pm 1.574$ $1.829\pm 0.009$ -
CRSVP-3 $0.899\pm 0.003$ $1.946\pm 0.025$ $1.691\pm 0.009$ $0.103\pm 0.014$ $0.899\pm 0.009$ $2.184\pm 0.063$ $1.394\pm 0.018$ $0.114\pm 0.009$
Table C4: Additional results for Caltech-101 and DBpedia. Coverage, efficiency, representation complexity and Pearson correlation values for the following unrestricted set-valued predictors: LAC, NPS, APS, and restricted set-valued predictors: NCRSVP, CRSVP, NCRSVP-$r$, and CRSVP-$r$. The confidence level is set to 90%, and calibration and test sets are resampled 10 times.
Dataset Caltech-101 DBpedia
Alg. Cov. Size Repr. comp. Corr. Cov. Size Repr. comp. Corr.
LAC $0.900\pm 0.007$ $0.920\pm 0.008$ $1.000\pm 0.000$ $0.886\pm 0.021$ $0.899\pm 0.001$ $0.931\pm 0.002$ $1.000\pm 0.000$ $0.814\pm 0.005$
NPS $1.000\pm 0.000$ $96.54\pm 0.065$ $1.454\pm 0.072$ - $1.000\pm 0.000$ $59.17\pm 0.272$ $53.18\pm 0.257$ $0.010\pm 0.003$
APS $0.900\pm 0.006$ $1.165\pm 0.015$ $1.251\pm 0.015$ $0.184\pm 0.013$ $0.901\pm 0.002$ $1.33\pm 0.287 11$$1.90\pm 0.277$ $0.129\pm 0.001$
NCRSVP $1.000\pm 0.000$ $96.25\pm 0.049$ $1.000\pm 0.000$ - $0.999\pm 0.000$ $162.9\pm 0.492$ $1.000\pm 0.000$ $0.019\pm 0.001$
CRSVP $0.901\pm 0.005$ $4.400\pm 0.254$ $1.000\pm 0.000$ $0.062\pm 0.009$ $0.901\pm 0.002$ $26.62\pm 0.417$ $1.000\pm 0.000$ $0.087\pm 0.002$
NCRSVP-1 $1.000\pm 0.000$ $70.36\pm 0.550$ $1.000\pm 0.000$ - $1.000\pm 0.000$ $183.2\pm 0.380$ $1.000\pm 0.000$ $0.023\pm 0.003$
CRSVP-1 $0.900\pm 0.005$ $4.173\pm 0.241$ $1.000\pm 0.000$ $0.079\pm 0.006$ $0.901\pm 0.002$ $25.59\pm 0.386$ $1.000\pm 0.000$ $0.119\pm 0.002$
NCRSVP-2 $1.000\pm 0.000$ $66.39\pm 0.528$ $1.115\pm 0.004$ - $0.999\pm 0.000$ $150.6\pm 0.418$ $1.351\pm 0.002$ $0.013\pm 0.001$
CRSVP-2 $0.900\pm 0.007$ $2.381\pm 0.110$ $1.085\pm 0.004$ $0.054\pm 0.014$ $0.901\pm 0.002$ $9.34\pm 0.484 1.1$$191\pm 0.002$ $0.107\pm 0.001$
NCRSVP-3 $1.000\pm 0.000$ $63.61\pm 0.517$ $1.306\pm 0.008$ - $0.999\pm 0.000$ $127.9\pm 0.501$ $1.985\pm 0.004$ $0.004\pm 0.001$
CRSVP-3 $0.901\pm 0.006$ $784\pm 0.086 1.1$$133\pm 0.007$ $0.066\pm 0.011$ $0.901\pm 0.001$ $17.71\pm 0.423$ $1.380\pm 0.005$ $0.113\pm 0.001$
Table C5: Additional results for Caltech-256 and PlantCLEF 2015. Coverage, efficiency, representation complexity and Pearson correlation values for the following unrestricted set-valued predictors: LAC, NPS, APS, and restricted set-valued predictors: NCRSVP, CRSVP, NCRSVP-$r$, and CRSVP-$r$. The confidence level is set to 90%, and calibration and test sets are resampled 10 times.
Dataset Caltech-256 PlantCLEF 2015
Alg. Cov. Size Repr. comp. Corr. Cov. Size Repr. comp. Corr.
LAC $0.900\pm 0.004$ $1.931\pm 0.040$ $1.926\pm 0.040$ $0.29\pm 0.006$ $0.899\pm 0.003$ $25.50\pm 0.450$ $24.33\pm 0.419$ $0.20\pm 0.009$
NPS $0.999\pm 0.000$ $62.25\pm 1.330$ $50.12\pm 0.996$ $0.01\pm 0.009$ $0.970\pm 0.001$ $123.6\pm 2.026$ $104.4\pm 1.560$ $0.003\pm 0.006$
APS $0.901\pm 0.004$ $3.640\pm 0.099$ $3.680\pm 0.098$ $0.05\pm 0.008$ $0.899\pm 0.003$ $44.12\pm 0.994$ $40.19\pm 0.836$ $0.01\pm 0.010$
NCRSVP $0.999\pm 0.000$ $208.4\pm 1.780$ $1.000\pm 0.000$ - $1.000\pm 0.000$ $998.9\pm 0.224$ $1.000\pm 0.000$ -
CRSVP $0.900\pm 0.003$ $44.69\pm 1.252$ $1.000\pm 0.000$ $0.100\pm 0.007$ $0.900\pm 0.002$ $520.9\pm 4.745$ $1.000\pm 0.000$ $0.340\pm 0.004$
NCRSVP-1 $0.999\pm 0.000$ $215.5\pm 1.885$ $1.000\pm 0.000$ - $1.000\pm 0.000$ $998.9\pm 0.176$ $1.000\pm 0.000$ -
CRSVP-1 $0.901\pm 0.003$ $44.24\pm 1.297$ $1.000\pm 0.000$ $0.139\pm 0.004$ $0.900\pm 0.003$ $534.5\pm 4.492$ $1.000\pm 0.000$ $0.356\pm 0.005$
NCRSVP-2 $0.999\pm 0.000$ $179.0\pm 2.035$ $1.242\pm 0.004$ - $1.000\pm 0.000$ $997.0\pm 0.443$ $1.002\pm 0.001$ -
CRSVP-2 $0.901\pm 0.003$ $27.63\pm 0.898$ $1.270\pm 0.006$ $0.073\pm 0.007$ $0.898\pm 0.003$ $442.1\pm 4.337$ $1.290\pm 0.004$ $0.295\pm 0.004$
NCRSVP-3 $1.000\pm 0.000$ $163.6\pm 2.070$ $1.632\pm 0.005$ - $1.000\pm 0.000$ $995.3\pm 0.492$ $1.006\pm 0.001$ -
CRSVP-3 $0.901\pm 0.003$ $20.30\pm 0.830$ $1.498\pm 0.009$ $0.045\pm 0.009$ $0.899\pm 0.004$ $389.7\pm 5.898$ $1.632\pm 0.010$ $0.262\pm 0.004$
Figure C4: Representation complexity versus set size for the following unrestricted set-valued predictors: LAC and APS, and restricted set-valued predictors: CRSVP and CRSVP-$r$, for all datasets. The confidence level is set to 90%, and calibration and test sets are resampled 10 times.